### v2

# Localized frames without inequalities

Abstract.
We consider countable families of vectors in a separable Hilbert space, which are mutually localized with respect to a fixed localized Riesz basis. We prove the equivalence of the frame property and nine conditions that do not involve any inequalities. This is done by studying the properties of their frame-related operators on the co-orbit spaces generated by the reference Riesz basis. We apply our main result to the setting of shift-invariant spaces and obtain new conditions for stable sets of sampling.
Key words and phrases:
localized frames, co-orbit spaces, R-dual sequences, invertibility of frame-related operators
2020 Mathematics Subject Classification:
42C15, 46B15, 47B37, 42B35, 46B45
1. Introduction
A frame is a natural generalization of the notion of orthonormal bases for a Hilbert space that allows for redundant representations, making them particularly suited for applications, for example, in signal processing. A frame is defined as a countable family of vectors $(\varphi_{k})_{k\in X}$ for which the famous chain of inequalities
(1) $A\,\|f\|^{2}\leq\sum_{k\in X}|\langle f,\varphi_{k}\rangle|^{2}\leq B\,\|f\|^{2},\qquad f\in\mathcal{H},$
is satisfied. In the article [21], Gröchenig gave fourteen equivalent conditions without inequalities for a Gabor system over a lattice with non-zero window in the modulation space $M^{1}({\mathbb{R}}^{d})$ being a frame. The main result of [21] is that any such Gabor system ${\mathcal{G}}={\mathcal{G}}(g,\Lambda)$ is a frame for $L^{2}(\mathbb{R}^{d})$ if and only if the associated frame related operators satisfy certain mapping properties, for example, if the analysis operator $C_{{\mathcal{G}}}:M^{\infty}(\mathbb{R}^{d})\to\ell^{\infty}(\Lambda)$ is injective. The proof of the latter relies on deep results of Gabor systems over lattices, such as the Ron-Shen duality principle [31], Wiener’s Lemma for twisted convolutions [23], and a result on linear independence of time-frequency shifts [21].
In many settings, however, frame theoretic results are ultimately not the consequence of an underlying group structure, but follow from localization properties of the frame, see, e.g., kernel theorems (compare [6, 14] with [10]), invertibility results of Gabor frame operators on modulation spaces [19, 23, 25], or a partial generalization [25, Theorem 5.1] of [21, Theorem 3.1]. A frame $(\psi_{k})_{k\in X}$ is said to be localized, if its Gram matrix is contained in a spectral matrix algebra (see [15] or Definition 2). To a localized frame $\psi$ one can then associate a family of Banach spaces, the so-called co-orbit spaces ${\mathcal{H}}^{p}(\psi)\ (1\leq p\leq\infty)$ which are defined in terms of summability conditions on the analysis coefficients $(\langle f,\psi_{k}\rangle)_{k\in X}$, see Section 2.3 for further details.
Returning to Gabor analysis, we note that the property $g\in M^{1}(\mathbb{R}^{d})$ of the window function can be seen as a localization assumption. Indeed, in the case $\Lambda=\mathbb{Z}^{d}$, we note that ${\mathcal{G}}(g,\mathbb{Z}^{d})$ is localized with respect to the unweighted Baskakov-Gohberg-Sjöstrand algebra $\mathcal{C}$ (see Example 4 (3)), and in the case of more irregular sets $\Lambda$, similar (more general) properties have been shown [25]. Furthermore, if ${\mathcal{G}}$ is a frame, the modulation spaces $M^{p}({\mathbb{R}}^{d})$ coincide with the co-orbit spaces $\mathcal{H}^{p}({\mathcal{G}})$ ($1\leq p\leq\infty$) generated from ${\mathcal{G}}$ in this case [25, Theorem 3.2]. Naturally, these observations lead to the question, whether abstract localized frames can be characterized in a similar fashion as in [21], namely via properties of their associated frame-related operators on their respective co-orbit spaces.
In this article we give several equivalent conditions for a suitably localized countable family $\psi=(\psi_{k})_{k\in X}$ of vectors in some abstract separable Hilbert space $\mathcal{H}$ being a frame. Let $C_{\psi}$, $D_{\psi}$ and $S_{\psi}$ denote the analysis, synthesis and frame operator respectively, and set $\text{Ran}_{p}(C_{\psi}):=C_{\psi}({\mathcal{H}}^{p}(\varphi))$ and $\text{Ran}_{p}(D_{\psi}):=D_{\psi}(\ell^{p}(X))$, see Section 2 for the detailed definitions. The formulation of our main result is as follows:
Theorem 1.
Let ${\mathcal{A}}$ be a spectral algebra and $\varphi=(\varphi_{k})_{k\in X}$ be an intrinsically ${\mathcal{A}}$-localized Riesz basis in $\mathcal{H}$. Let $\psi=(\psi_{k})_{k\in X}$ be a family in $\mathcal{H}$, which is mutually ${\mathcal{A}}$-localized with respect to $\varphi$, and let $\omega=(\omega_{k})_{k\in X}$ be given by
$\omega_{k}=\sum_{l\in X}\langle\psi_{l},\varphi_{k}\rangle S_{\varphi}^{-{1}/{2}}\varphi_{l},\qquad k\in X.$
Then the following are equivalent.
  1. (1)
    $\psi$ is a frame for $\mathcal{H}$.
  2. (2)
    $S_{\psi}$ is invertible on $\mathcal{H}^{1}(\varphi)$.
  3. (3)
    $S_{\psi}$ is invertible on $\mathcal{H}^{\infty}(\varphi)$.
  4. (4)
    $C_{\psi}:\mathcal{H}^{\infty}(\varphi)\rightarrow\ell^{\infty}(X)$ is injective and $\emph{Ran}_{\infty}(C_{\psi})$ as well as $\emph{Ran}_{\infty}(D_{\psi})$ are closed.
  5. (5)
    $D_{\psi}:\ell^{1}(X)\rightarrow\mathcal{H}^{1}(\varphi)$ is surjective and $\emph{Ran}_{1}(C_{\psi})$ is closed.
  6. (6)
    $D_{\omega}:\ell^{\infty}(X)\rightarrow\mathcal{H}^{\infty}(\varphi)$ is injective and $\emph{Ran}_{\infty}(C_{\omega})$ as well as $\emph{Ran}_{\infty}(D_{\omega})$ are closed.
  7. (7)
    $C_{\omega}:\mathcal{H}^{1}(\varphi)\rightarrow\ell^{1}(X)$ is surjective and $\emph{Ran}_{1}(D_{w})$ is closed.
  8. (8)
    $G_{\omega}$ is invertible on $\ell^{1}(X)$.
  9. (9)
    $G_{\omega}$ is invertible on $\ell^{\infty}(X)$.
  10. (10)
    $\omega$ is a Riesz sequence in ${\mathcal{H}}$.
In particular, if any of the above conditions is satisfied, then $\mathcal{H}^{p}(\psi)=\mathcal{H}^{p}(\varphi)$ with equivalent norms ($1\leq p\leq\infty$).
Naturally, the price for working in an abstract Hilbert space is that less structure and less tools are available. This forces us to impose additional assumptions in the statements of Theorem 1, when compared with [21].
First, while the modulation spaces $M^{p}({\mathbb{R}}^{d})$ can be intrinsically defined via the short-time Fourier transform associated with the omnipresent Gaussian (or any other Schwartz function) [18], the definition and many relevant properties of the co-orbit spaces $\mathcal{H}^{p}(\varphi)$ depend on the frame property of some explicitly given localized frame $\varphi$. This is the reason why we relate the family $\psi$ to an intrinsically localized reference frame $\varphi$ generating the co-orbit spaces $\mathcal{H}^{p}(\varphi)$, instead of working with the co-orbit spaces $\mathcal{H}^{p}(\psi)$ directly.
Second, to our knowledge, there is currently no canonical generalization of the Ron-Shen duality principle to abstract frames in the literature. The Ron-Shen duality principle connects any Gabor frame on a lattice with its (dual) Gabor Riesz sequence generated by the same window on the adjoint lattice. In Theorem 1 we consider the family $\omega$, which, in the language of [33], is a so-called R-dual of type III of $\psi$, and serves as a substitute of the aforementioned Ron-Shen dual. In general, however, an R-dual associated with a Gabor frame on a lattice does not coincide with its Ron-Shen dual, see [11, 33] for a discussion of this matter. In order to guarantee that the R-dual $\omega$ is a Riesz sequence if and only if $\psi$ is a frame, we additionally have to assume that the reference system $\varphi$ is in fact a Riesz basis and not only a frame. Thus, our localization assumption $\psi\sim_{\mathcal{A}}\varphi$ corresponds to the original localization concept from [19] in the case ${\mathcal{A}}={\mathcal{J}}_{s}$ (see Example 4 (1)), which was followed by several other publications on this topic [1, 3, 4, 5, 15, 20]. At the same time, however, it might be very difficult, if not even impossible for a Gabor frame to be localized with respect to a localized Riesz basis (see also [19, Section 5] for a discussion of this issue).
Finally, several closed range conditions appear in our setting but not in [21, Theorem 3.1]. We show in Example 28 and Remark 29, that each of the closed range conditions appearing in Theorem 1 is in fact necessary in our setting, since otherwise, the statement is simply not true. For Gabor frames, however, these conditions are automatically satisfied, see Example 27.
By the previous observations, Theorem 1 does not directly generalize [21, Theorem 3.1]. However, [21] served as inspiration for this article, and multiple equivalent conditions of [21, Theorem 3.1] can be translated directly into conditions in Theorem 1. We refer to Section 5 for a more detailed comparison.
This article is structured as follows. In Section 2 we fix our notation and summarize some relevant concepts and results from the literature. In Section 3 we prove some auxiliary results, which we utilize in Section 4, where we prove our main result. In Section 5 we compare Theorem 1 to [21, Theorem 3.1], which initiated the research on the content of this article. Finally, Section 6 is devoted to applying our main result in the setting of shift-invariant spaces.
2. Preliminaries and Notation
[…]
$\big{\langle}(c_{k})_{k\in X},(d_{k})_{k\in X}\big{\rangle}_{\ell^{p}\times\ell^{q}}=\sum_{k\in X}c_{k}\overline{d_{k}}$
in order to be compatible with the Hilbert space duality on $\ell^{2}(X)$.
2.1. Basic Frame Theory
A countable family $\varphi:=(\varphi_{k})_{k\in X}$ of vectors in ${\mathcal{H}}$ is called a frame, if there exist constants $A,B>0$ such that
(2) $A\,\|f\|^{2}\leq\sum_{k\in X}|\langle f,\varphi_{k}\rangle|^{2}\leq B\,\|f\|^{2},\qquad f\in\mathcal{H}.$
For a comprehensive introduction to frames in Hilbert space we refer the reader to [12]. If the upper, but not necessarily the lower inequality in (2) is satisfied for some prescribed $B>0$ and all $f\in\mathcal{H}$, then $\varphi$ is called a Bessel sequence. Whenever $\varphi$ is a Bessel sequence, both the analysis operator, defined as
$C_{\varphi}:{\mathcal{H}}\rightarrow\ell^{2}(X),\qquad C_{\varphi}f=\big{(}\langle f,\varphi_{k}\rangle\big{)}_{k\in X},$
and the synthesis operator, defined as
$D_{\varphi}:\ell^{2}(X)\rightarrow{\mathcal{H}},\qquad D_{\varphi}c=\sum_{k\in X}c_{k}\varphi_{k},$
are bounded and adjoint to one another, where the latter series converges unconditionally in $\mathcal{H}$. A combination of the analysis and the synthesis operator yields the frame operator
$S_{\varphi}=D_{\varphi}C_{\varphi}:{\mathcal{H}}\rightarrow{\mathcal{H}},\qquad S_{\varphi}f=\sum_{k\in X}\langle f,\varphi_{k}\rangle\varphi_{k}.$
If $\varphi$ is a frame, then $S_{\varphi}$ is bounded, positive, self-adjoint and invertible and the family $\widetilde{\varphi}=\big{(}\widetilde{\varphi}_{k}\big{)}_{k\in X}:=\big{(}S_{\varphi}^{-1}\varphi_{k}\big{)}_{k\in X}$ is another frame, called the canonical dual frame, that satisfies
(3) $f=D_{\varphi}C_{\widetilde{\varphi}}f=\sum_{k\in X}\langle f,\widetilde{\varphi}_{k}\rangle\varphi_{k}=D_{\widetilde{\varphi}}C_{\varphi}f=\sum_{k\in X}\langle f,\varphi_{k}\rangle\widetilde{\varphi}_{k},\qquad f\in\mathcal{H}.$
More generally, if $\varphi^{d}=(\varphi^{d})_{k\in X}$ is another family in $\mathcal{H}$ such that $D_{\varphi}C_{\varphi^{d}}=D_{\varphi^{d}}C_{\varphi}=\mathcal{I}_{\mathcal{H}}$, then $\varphi^{d}$ is called a dual frame of $\varphi$. If $\psi,\varphi$ are two countable families of vectors in ${\mathcal{H}}$ indexed by $X$, then we define the cross Gram matrix as
(4) $G_{\psi,\varphi}=\big{[}\langle\varphi_{l},\psi_{k}\rangle\big{]}_{k,l\in X}.$
If $\psi=\varphi$ we write $G_{\varphi}=G_{\varphi,\varphi}$. In general, a (cross) Gram matrix as in (4) might not be bounded on $\ell^{2}(X)$, but in case that $\psi$ and $\varphi$ both are Bessel sequences, then $G_{\psi,\varphi}$ defines a bounded operator on $\ell^{2}(X)$ and
$G_{\psi,\varphi}=C_{\psi}D_{\varphi}$
and $G_{\psi,\varphi}^{\ast}=G_{\varphi,\psi}$. In fact, a countable family $\varphi$ in $\mathcal{H}$ is a Bessel sequence if and only if its associated Gram matrix $G_{\varphi}$ defines a bounded operator on $\ell^{2}(X)$ [7]. A family $\varphi:=(\varphi_{k})_{k\in X}\subset{\mathcal{H}}$ is called a Riesz sequence for ${\mathcal{H}}$, if there exist constants $A,B>0$ such that
$A\|c\|_{\ell^{2}(X)}^{2}\leq\left\|\sum_{k\in X}c_{k}\varphi_{k}\right\|^{2}\leq B\|c\|_{\ell^{2}(X)}^{2},\qquad c\in\ell^{2}(X).$
Equivalently, $\varphi$ is a Riesz sequence if and only if its associated Gram matrix $G_{\varphi}$ defines a bounded and invertible operator on $\ell^{2}(X)$ [7]. A Riesz sequence $\varphi$ is called a Riesz basis, if it is also complete, that is, $\overline{\text{span}}(\varphi_{k})_{k\in X}={\mathcal{H}}$. In particular, a Riesz basis is a frame for which both $G_{\varphi}\in{\mathcal{B}}(\ell^{2}(X))$ and $S_{\varphi}\in{\mathcal{B}}(\mathcal{H})$ are invertible. For any Riesz basis $\varphi$ the family $(S_{\varphi}^{-1/2}\varphi_{k})_{k\in X}$ constitutes an orthonormal basis.
2.2. The R-dual sequence
We will consider the following notion of duality for abstract sequences in ${\mathcal{H}}$. Let $(e_{k})_{k\in X}$ and $(h_{k})_{k\in X}$ be two orthonormal bases for ${\mathcal{H}}$, and let $(\varphi_{k})_{k\in X}$ be a family of vectors such that $\sum_{n\in X}|\langle\varphi_{n},e_{k}\rangle|^{2}<\infty$ for every $k\in X$. Then
$\omega_{k}=\sum_{n\in X}\langle\varphi_{n},e_{k}\rangle h_{n}$
converges in $\mathcal{H}$ for each $k\in X$ and the family $\omega=(\omega_{k})_{k\in X}$ is called an R-dual or Riesz-dual sequence for $\varphi$. It was shown in [11, Propositions 10,11] that $\varphi$ is a frame for ${\mathcal{H}}$ if and only if $\omega$ is a Riesz sequence.
2.3. Localized Frames
As mentioned in the previous section, whenever $\varphi$ is a Bessel sequence, its associated Gram matrix $G_{\varphi}=[\langle\varphi_{l},\varphi_{k}\rangle]_{k,l\in X}$ defines a bounded operator on $\ell^{2}(X)$. Boundedness on $\ell^{2}(X)$ is equivalent to some form of off-diagonal decay of the matrix entries $\langle\varphi_{l},\varphi_{k}\rangle$, as described, e.g., in [30, Theorem 4.16]. The idea of localization is to impose stronger off-diagonal decay conditions on a given (cross) Gram matrix. Off-diagonal decay of matrices is often described by Banach matrix algebras, which motivated the authors in [15] to consider the following type of matrix algebras.
Definition 2.
A unital Banach *-algebra $\mathcal{A}$ of $\mathbb{C}$-valued matrices $A=[A_{k,l}]_{k,l\in X}$ is called a solid spectral matrix algebra [15, Section 2.3], or simply spectral algebra, if
  1. (A0)
    ${\mathcal{A}}\subset\bigcap_{1\leq p\leq\infty}{\mathcal{B}}(\ell^{p}(X))$, i.e., each $A\in\mathcal{A}$ defines a bounded operator on $\ell^{p}(X)$ for each $1\leq p\leq\infty$.
  2. (A1)
    ${\mathcal{A}}\subset{\mathcal{B}}(\ell^{2}(X))$, i.e., each $A\in\mathcal{A}$ defines a bounded operator on $\ell^{2}(X)$.
  3. (A2)
    $\mathcal{A}$ is inverse-closed in $\mathcal{B}(\ell^{2}(X))$, i.e., if $A\in\mathcal{A}$ is invertible in $\mathcal{B}(\ell^{2}(X))$, then its bounded inverse $A^{-1}\in{\mathcal{B}}(\ell^{2}(X))$ is contained in $\mathcal{A}$ as well.
  4. (A3)
    $\mathcal{A}$ is solid, i.e., if $A=[A_{k,l}]_{k,l\in X}\in\mathcal{A}$ and $|B_{k,l}|\leq|A_{k,l}|$ for all $k,l\in X$, then $B=[B_{k,l}]_{k,l\in X}\in\mathcal{A}$ and $\|B\|_{\mathcal{A}}\leq\|A\|_{\mathcal{A}}$.
Remark 3.
Condition (A2) is equivalent to spectral invariance of ${\mathcal{A}}$ in ${\mathcal{B}}(\ell^{2}(X))$, that is, $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$ for all $A\in{\mathcal{A}}$ [22, Lemma 2.4], where $\sigma_{{\mathcal{A}}}(A)$ and $\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$ denote the spectrum of $A$ in the Banach algebras ${\mathcal{A}}$ and ${\mathcal{B}}(\ell^{2}(X))$ respectively. As a consequence, the holomorphic (Riesz) functional calculi (see e.g. [13, Chap. VII]) of ${\mathcal{A}}$ and ${\mathcal{B}}(\ell^{2}(X))$ coincide. More precisely, if $A\in{\mathcal{A}}$, $U\subseteq\mathbb{C}$ is an open neighborhood of (a connected component of) $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$, $\Gamma\subseteq U$ is a contour of $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$, and if $z\mapsto f(z)$ is holomorphic on $U$, then
[…]
defines the same element in ${\mathcal{A}}$ and ${\mathcal{B}}(\ell^{2}(X))$.
Strictly speaking, the authors in [15] considered algebras satisfying properties (A1--A3) without including property (A0) in their definition. In the language of [15, Definition 2], property (A0) can be rephrased to saying that the constant weight $1$ is ${\mathcal{A}}$-admissible with critical index $p_{0}=1$. Here we assume for simplicity, that a spectral algebra automatically satisfies (A0) as well, since many of the classical examples of spectral algebras do so anyway, see below.
Example 4.
Let $X\subset\mathbb{R}^{d}$ be a relatively separated set. Then the following classes of matrices are examples of spectral algebras in the sense of Definition 2, see also [29] for their ${\mathcal{B}}(\mathcal{H})$-valued analogs.
(1) For every $s>d$, the Jaffard algebra ${\mathcal{J}}_{s}={\mathcal{J}}_{s}(X)$ of matrices $A=[A_{k,l}]_{k,l\in X}$ for which
$\sup_{k,l\in X}|A_{k,l}|(1+|k-l|)^{s}<\infty$
is a spectral algebra [27].
(2) For every continuous function $\nu:\mathbb{R}^{d}\to[0,\infty)$ which
  • (2a)
[…]
  • (2b)
    satisfies the GRS-condition (Gelfand-Raikov-Shilov condition [16])
    $\lim_{n\rightarrow\infty}(\nu(nz))^{\frac{1}{n}}=1,\qquad z\in\mathbb{R}^{d},$
  • (2c)
    satisfies the weak growth condition
    $\nu(x)\geq C(1+|x|)^{\delta}\qquad\text{for some }\delta\in(0,1],C>0,$
the weighted Schur algebra ${\mathcal{S}}^{1}_{\nu}={\mathcal{S}}^{1}_{\nu}(X)$ of matrices $A=[A_{k,l}]_{k,l\in X}$ for which
$\max\left\{\sup_{k\in X}\sum_{l\in X}|A_{k,l}|\nu(k-l)\,,\,\sup_{l\in X}\sum_{k\in X}|A_{k,l}|\nu(k-l)\right\}<\infty$
is a spectral algebra [24, 34].
(3) For every continuous function $\nu:\mathbb{R}^{d}\to[0,\infty)$ which
  • (3a)
    is submultiplicative, i.e., $\nu(x+y)\leq\nu(x)\nu(y)$ for every $x,y\in\mathbb{R}^{d}$,
  • (3b)
    is symmetric, i.e., $\nu(-x)=\nu(x)$ for every $x\in\mathbb{R}^{d}$,
  • (3c)
    satisfies the GRS-condition,
the weighted Baskakov-Gohberg-Sjöstrand algebra $\mathcal{C}_{\nu}=\mathcal{C}_{\nu}(\mathbb{Z}^{d})$ of matrices $A=[A_{k,l}]_{k,l\in\mathbb{Z}^{d}}$ for which
$\sum_{l\in\mathbb{Z}^{d}}\sup_{k\in\mathbb{Z}^{d}}|A_{k,k-l}|\nu(l)<\infty$
is a spectral algebra [8, 17, 32].
(4) Anisotropic versions of each of the aforementioned algebras are further examples of spectral algebras [29, Section 6].
For a scalar matrix $A=[A_{k,l}]_{k,l\in X}$, we define the complex conjugate $\overline{A}$ of $A$ and the absolute value $|A|$ of $A$ pointwise, i.e., by $\overline{A}:=\big{[}\,\overline{A_{k,l}}\,\big{]}_{k,l\in X}$ and $|A|:=\big{[}\,|A_{k,l}|\,\big{]}_{k,l\in X}$. We explicitly note that by the solidity condition (A3), it holds
$A\in{\mathcal{A}}\qquad\Longleftrightarrow\qquad\overline{A}\in{\mathcal{A}}\qquad\Longleftrightarrow\qquad|A|\in{\mathcal{A}}.$
We will use this fact several times without further reference.
Definition 5.
[…]
Note that by involution in ${\mathcal{A}}$ it holds that $\psi\sim_{{\mathcal{A}}}\varphi$ if and only if $\varphi\sim_{{\mathcal{A}}}\psi$. We will use this fact without further reference.
One of the main results in [15] states that localization of a frame is preserved under canonical duality. More precisely:
Theorem 6.
[15, Theorem 3.6] Let $\varphi$ be a frame in $\mathcal{H}$. If $\varphi\sim_{{\mathcal{A}}}\varphi$ then also $\widetilde{\varphi}\sim_{{\mathcal{A}}}\widetilde{\varphi}$ and $\varphi\sim_{{\mathcal{A}}}\widetilde{\varphi}$.
The latter theorem immediately implies the following simple lemma, which we will also implicitly apply several times without further notice. For completeness we present the simple proof.
Lemma 7.
Let $\varphi=(\varphi_{k})_{k\in X}$ an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ and $\psi=(\psi_{k})_{k\in X}$ be any countable family in $\mathcal{H}$ such that $\psi\sim_{{\mathcal{A}}}\varphi$. Then it holds $\psi\sim_{{\mathcal{A}}}\widetilde{\varphi}$ and $\psi\sim_{{\mathcal{A}}}\psi$.
Proof.
By assumption, $G_{\psi,\varphi}\in{\mathcal{A}}$ and by Theorem 6, $G_{\widetilde{\varphi}}\in{\mathcal{A}}$. Thus, by frame reconstruction and the algebra property of ${\mathcal{A}}$, $G_{\psi,\widetilde{\varphi}}=G_{\psi,\varphi}G_{\widetilde{\varphi}}\in{\mathcal{A}}$. Similarly, the latter implies that $G_{\psi}=G_{\psi,\varphi}G_{\widetilde{\varphi},\psi}\in{\mathcal{A}}$. ∎
A remarkable feature of a localized frame $\varphi$ in $\mathcal{H}$ is, that the reconstruction formula (3) does not only converge in the underlying Hilbert space $\mathcal{H}$, but also in a whole chain of associated Banach spaces defined below.
Definition 8.
[15, Definition 5] Let $\mathcal{A}$ be a spectral algebra, $\varphi$ be a frame for $\mathcal{H}$, and $\varphi^{d}$ be some dual frame of $\varphi$ such that $\varphi\sim_{{\mathcal{A}}}\varphi^{d}$. Let
(5) $\begin{split}{\mathcal{H}}^{00}(\varphi):=D_{\varphi}(\ell^{00}(X))=\left\{\sum_{k\in X}c_{k}\varphi_{k}:\ c\in\ell^{00}(X)\right\}\end{split}$
be the space of finite linear combinations of elements from $\varphi$. For $1\leq p<\infty,$ the co-orbit space ${\mathcal{H}}^{p}(\varphi^{d},\varphi)$ is defined as the norm completion of ${\mathcal{H}}^{00}(\varphi)$ with respect to the norm $\|f\|_{{\mathcal{H}}^{p}(\varphi^{d},\varphi)}:=\big{\|}C_{\varphi^{d}}f\big{\|}_{\ell^{p}(X)}$.
Under the above assumptions, the spaces are well-defined non-trivial Banach spaces with ${\mathcal{H}}^{p}(\varphi^{d},\varphi)\subseteq{\mathcal{H}}^{q}(\varphi^{d},\varphi)$ for all $1\leq p,q<\infty$ and ${\mathcal{H}}^{2}(\varphi^{d},\varphi)=\mathcal{H}$. Furthermore, the analysis operator $C_{\widetilde{\varphi}}:\mathcal{H}^{00}(\varphi)\to\ell^{00}(X)$ extends via density to an isometry $C_{\widetilde{\varphi}}:\mathcal{H}^{p}(\varphi)\to\ell^{p}(X)$ ($1\leq p<\infty$).
Remark 9.
The definition of ${\mathcal{H}}^{p}(\varphi^{d},\varphi)$ seems to depend heavily on the dual frame pair $(\varphi,\varphi^{d})$. However, if $(\psi,\psi^{d})$ is another pair of dual frames satisfying $\psi\sim_{{\mathcal{A}}}\psi^{d}$, and if $\psi^{d}\sim_{{\mathcal{A}}}\varphi$ and $\varphi^{d}\sim_{{\mathcal{A}}}\psi$, then $\mathcal{H}^{p}(\varphi^{d},\varphi)=\mathcal{H}^{p}(\psi^{d},\psi)$ with equivalent norms [28, Theorem 3.25].
If $\varphi$ is an intrinsically ${\mathcal{A}}$-localized frame, then $\varphi\sim_{{\mathcal{A}}}\widetilde{\varphi}$ and $\widetilde{\varphi}\sim_{{\mathcal{A}}}\widetilde{\varphi}$ by Theorem 6. In particular, due to the above remark, it holds $\mathcal{H}^{p}(\widetilde{\varphi},\varphi)=\mathcal{H}^{p}(\varphi,\widetilde{\varphi})$ with equivalent norms and we write $\mathcal{H}^{p}(\varphi):=\mathcal{H}^{p}(\widetilde{\varphi},\varphi)$ in this case ($1\leq p<\infty$).
The definition of the space $\mathcal{H}^{\infty}(\varphi)$ associated to an ${\mathcal{A}}$-localized frame $\varphi$ in $\mathcal{H}$ is more subtle. One could define it as the anti-linear topological dual space of $\mathcal{H}^{1}(\varphi)$ (see Theorem 15), or as the completion of $\mathcal{H}$ with respect to the Hausdorff locally convex $\sigma(\mathcal{H},\mathcal{H}^{00}(\widetilde{\varphi}))$-topology as done in [15, 19]. In this article, we work with the definition from [5] (see also [28]) given below.
Definition 10.
[5, 28] Let $(f_{n})_{n\in{\mathbb{N}}}$ and $(g_{n})_{n\in{\mathbb{N}}}$ be sequences in $\mathcal{H}$ and $\varphi$ be a frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then we call $(f_{n})_{n\in{\mathbb{N}}}$ and $(g_{n})_{n\in{\mathbb{N}}}$ equivalent, in short $(f_{n})_{n\in{\mathbb{N}}}\sim_{\varphi}(g_{n})_{n\in{\mathbb{N}}}$, if
$\lim_{n\rightarrow\infty}\langle f_{n}-g_{n},\widetilde{\varphi}_{k}\rangle=0\qquad(\forall k\in X)\qquad\text{and}\qquad\sup_{n\in\mathbb{N}}\|C_{\widetilde{\varphi}}(f_{n}-g_{n})\|_{\ell^{\infty}(X)}<\infty.$
Clearly the relation $\sim_{\varphi}$ defines an equivalence relation on the space of all sequences in $\mathcal{H}$.
Definition 11.
[5, 28] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then $\mathcal{H}^{\infty}(\varphi)$ is defined to be the space of all equivalence classes $f=[(f_{n})_{n\in{\mathbb{N}}}]_{\sim_{\varphi}}$ of sequences $(f_{n})_{n\in{\mathbb{N}}}$ in $\mathcal{H}$ satisfying
  • (i)
    $\lim_{n\rightarrow\infty}\langle f_{n},\widetilde{\varphi}_{k}\rangle=:\langle f,\widetilde{\varphi}_{k}\rangle$ exists for each $k\in X$,
  • (ii)
    $\sup_{n\in\mathbb{N}}\|C_{\widetilde{\varphi}}f_{n}\|_{\ell^{\infty}(X)}<\infty.$
It is quickly verified that the defining properties (i) and (ii) do not depend on the choice of the representative sequence $(f_{n})_{n\in\mathbb{N}}$. Furthermore, condition (ii) implies that there is a positive constant $C$ such that
(6) $|\langle f_{n},\widetilde{\varphi}_{k}\rangle|\leq C\qquad\text{for each }k\in X\text{ and each }n\in\mathbb{N}.$
This implies that $|\langle f,\widetilde{\varphi}_{k}\rangle|\leq C$ for all $k\in X$. In particular,
(7) $\|f\|_{\mathcal{H}^{\infty}(\varphi)}:=\|C_{\widetilde{\varphi}}f\|_{\ell^{\infty}(X)}=\sup_{k\in X}\left|\lim_{n\rightarrow\infty}\langle f_{n},\widetilde{\varphi}_{k}\rangle\right|$
is well-defined and, in fact, defines a norm on $\mathcal{H}^{\infty}(\varphi)$ with respect to which $\mathcal{H}^{\infty}(\varphi)$ is complete. We explicitly emphasize that here $\langle f,\widetilde{\varphi}_{k}\rangle$ is not an inner product in $\mathcal{H}$ but (by abusing notation) the limit of the sequence $\big{(}\langle f_{n},\widetilde{\varphi}_{k}\rangle\big{)}_{n\in{\mathbb{N}}}$ (for each $k\in X$). Analogously, $C_{\widetilde{\varphi}}f=(\langle f,\widetilde{\varphi}_{k}\rangle)_{k\in X}$ is understood as a pointwise limit. In this sense, $C_{\widetilde{\varphi}}:\mathcal{H}^{\infty}(\varphi)\to\ell^{\infty}(X)$ defines an isometry.
Proposition 12.
[15, Theorem 3.8], [28, Corollary 3.27] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then $\mathcal{H}^{\infty}(\varphi)$ is a Banach space with respect to the norm $\|C_{\widetilde{\varphi}}\cdot\|_{\ell^{\infty}(X)}$, and $\|C_{\varphi}\cdot\|_{\ell^{\infty}(X)}$ is an equivalent norm.
The properties of the operators $D_{\varphi}$ and $S_{\varphi}$ on their respective domains are summarized below.
Proposition 13.
[15, Proposition 2.4],[28, Proposition 3.20] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then for each $1\leq p<\infty$, the synthesis operator
$D_{\varphi}:\ell^{00}(X)\longrightarrow\mathcal{H}^{00}(\varphi),$
(8) $(c_{l})_{l\in X}\mapsto\sum_{l\in X}c_{l}\varphi_{l}$
continuously extends to a bounded and surjective operator $D_{\varphi}:\ell^{p}(X)\to\mathcal{H}^{p}(\varphi)$ and the associated series (13) converges unconditionally in $\mathcal{H}^{p}(\varphi)$.
Similarly, if $F_{1}\subseteq F_{2}\subseteq\dots$ is a nested sequence of finite subsets of $X$ such that $\bigcup_{n=1}^{\infty}F_{n}=X$, then $f_{n}:=\sum_{l\in F_{n}}c_{l}\psi_{l}$ is contained in $\mathcal{H}$ for each $n\in\mathbb{N}$ and
(9) $c=(c_{l})_{l\in X}\mapsto D_{\varphi}c:=\big{[}(f_{n})_{n\in{\mathbb{N}}}\big{]}_{\sim_{\varphi}}=\left[\left(\sum_{l\in F_{n}}c_{l}\psi_{l}\right)_{n\in{\mathbb{N}}}\,\right]_{\sim_{\varphi}}$
defines a bounded and surjective operator $D_{\varphi}:\ell^{\infty}(X)\to\mathcal{H}^{\infty}(\varphi)$ and $D_{\varphi}c$ is independent of the choice of $\{F_{n}\}_{n=1}^{\infty}$, i.e. (9) converges unconditionally in the $\sigma(\mathcal{H},\mathcal{H}^{00}(\widetilde{\varphi}))$-topology.
Theorem 14.
[15, Theorem 3.8],[28, Corollary 3.27] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then for each $1\leq p\leq\infty$, the frame operator $S_{\varphi}=D_{\varphi}C_{\varphi}:\mathcal{H}^{p}(\varphi)\to\mathcal{H}^{p}(\varphi)$ is bounded and invertible.
We conclude this section with the following result on the anti-linear topological dual space of $\mathcal{H}^{p}(\varphi)$.
Theorem 15.
[5, Proposition 2] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then
$(\mathcal{H}^{p}(\varphi))^{\prime}=\mathcal{H}^{q}(\varphi)$
with equivalent norms for all $1\leq p<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, where the duality relation is given by
$\langle f,g\rangle_{\mathcal{H}^{p}\times\mathcal{H}^{q}}=\langle C_{\widetilde{\varphi}}f,C_{\varphi}g\rangle_{\ell^{p}\times\ell^{q}}=\sum_{k\in X}\left[C_{\widetilde{\varphi}}f\right]_{k}\overline{\left[C_{\varphi}g\right]_{k}}.$
3. Auxiliary Results
In this section we prepare the tools that will be important for the proof of our main theorem.
3.1. Frame-related Operators
First, we collect some preliminary results on properties of the frame-related operators associated with localized sequences.
Lemma 16.
Let $\varphi$ be a Riesz sequence in ${\mathcal{H}}$ that satisfies $\varphi\sim_{\mathcal{A}}\varphi$. Then $G_{\varphi}:\ell^{p}(X)\to\ell^{p}(X)$ is bounded and invertible for every $1\leq p\leq\infty$.
Proof.
By the localization assumption we have that $G_{\varphi}\in{\mathcal{A}}$. The Riesz sequence property implies that $G_{\varphi}$ is invertible on $\mathcal{B}(\ell^{2}(X))$, which implies by inverse closedness of ${\mathcal{A}}$ in $\mathcal{B}(\ell^{2}(X))$ that $(G_{\varphi})^{-1}\in{\mathcal{A}}$. Consequently, $G_{\varphi}$ is invertible on all $\ell^{p}(X)$-spaces ($1\leq p\leq\infty$). ∎
Corollary 17.
Let $\varphi$ be a Riesz basis for ${\mathcal{H}}$ that satisfies $\varphi\sim_{\mathcal{A}}\varphi$. Then $D_{\varphi}:\ell^{p}(X)\to{\mathcal{H}}^{p}(\varphi)$ is injective for every $1\leq p\leq\infty$.
Proof.
We may write $G_{\varphi}:\ell^{p}(X)\to\ell^{p}(X)$ as $G_{\varphi}=C_{\varphi}D_{\varphi}$ with $D_{\varphi}:\ell^{p}(X)\to{\mathcal{H}}^{p}(\varphi)$ being bounded (Proposition 13). The result then follows immediately from Lemma 16. ∎
The next two lemmata generalize [19, Proposition 8], where the case of the Jaffard algebra ${\mathcal{A}}=\mathcal{J}_{s}$ is treated. They are also similar to [5, Lemma 6 and 7], except for the frame assumption on $\psi$. In fact, our conclusion is also weaker due to the missing frame assumption, i.e., we cannot conclude here that $C_{\psi}$ has closed range, as explicitly demonstrated in Example 28 (iv).
Lemma 18.
Let $\varphi=(\varphi_{k})_{k\in X}$ be a frame in $\mathcal{H}$ and $\psi=(\psi_{k})_{k\in X}$ be a family in ${\mathcal{H}}$ that satisfies $\psi\sim_{\mathcal{A}}\varphi$. Then $C_{\psi}:{\mathcal{H}}^{p}(\varphi)\to\ell^{p}(X)$ is well-defined and bounded for each $1\leq p\leq\infty$. In particular, $\psi$ is a Bessel sequence in $\mathcal{H}$.
Proof.
First, let $1\leq p<\infty$. Let $f\in\mathcal{H}^{00}(\varphi)$ be arbitrary. Then $C_{\widetilde{\varphi}}f\in\ell^{p}(X)$ and
$\|C_{\psi}f\|_{\ell^{p}(X)}=\|G_{\psi,\varphi}C_{\widetilde{\varphi}}f\|_{\ell^{p}(X)}\leq\|G_{\psi,\varphi}\|_{{\mathcal{B}}(\ell^{p}(X))}\|f\|_{\mathcal{H}^{p}(\varphi)},$
where we used frame reconstruction in $\mathcal{H}^{00}(\varphi)\subseteq\mathcal{H}$ with respect to $\varphi$ and $\widetilde{\varphi}$, as well as property (A0) from Definition 2. Thus, by density, $C_{\psi}$ extends to a bounded operator ${\mathcal{H}}^{p}(\varphi)\to\ell^{p}(X)$ which we again denote by $C_{\psi}$.
Now let $p=\infty$. Let $f=\big{[}(f_{n})_{n\in{\mathbb{N}}}\big{]}_{\sim_{\varphi}}\in\mathcal{H}^{\infty}(\varphi)$ be arbitrary. Then, by definition, $\lim_{n\to\infty}\langle f_{n},\widetilde{\varphi}_{k}\rangle=:\langle f,\widetilde{\varphi}_{k}\rangle$ converges for every $k\in X$ and $\sup_{n\in\mathbb{N}}\|C_{\widetilde{\varphi}}f_{n}\|_{\ell^{\infty}(X)}$ $\leq C$ for some $C>0$, as well as $C_{\widetilde{\varphi}}f:=(\langle f,\widetilde{\varphi}_{k}\rangle)_{k\in X}\in\ell^{\infty}(X)$. This implies that
(10) $\sum_{l\in X}|\langle f_{n},\widetilde{\varphi}_{l}\rangle||\langle\varphi_{l},\psi_{k}\rangle|\leq C\sup_{k\in X}\sum_{l\in X}|\langle\varphi_{l},\psi_{k}\rangle|=C\|G_{\psi,\varphi}\|_{{\mathcal{B}}(\ell^{\infty}(X))}<\infty$
for each $k\in X$ and each $n\in\mathbb{N}$. Hence we may apply the dominated convergence theorem and obtain via frame reconstruction
$\lim_{n\to\infty}\langle f_{n},\psi_{k}\rangle=\lim_{n\to\infty}\sum_{l\in X}\langle f_{n},\widetilde{\varphi}_{l}\rangle\langle\varphi_{l},\psi_{k}\rangle=\sum_{l\in X}\langle f,\widetilde{\varphi}_{l}\rangle\langle\varphi_{l},\psi_{k}\rangle=[G_{\psi,\varphi}C_{\widetilde{\varphi}}f]_{k}$
for each $k\in X$. Consequently,
$\|C_{\psi}f\|_{\ell^{\infty}(X)}$ $=\left\|\left(\lim_{n\to\infty}\langle f_{n},\psi_{k}\rangle\right)_{k\in X}\right\|_{\ell^{\infty}(X)}$
$=\|G_{\psi,\varphi}C_{\widetilde{\varphi}}f\|_{\ell^{\infty}(X)}$
$\leq\|G_{\psi,\varphi}\|_{{\mathcal{B}}(\ell^{\infty}(X))}\|C_{\widetilde{\varphi}}f\|_{\ell^{\infty}(X)}$
$=\|G_{\psi,\varphi}\|_{{\mathcal{B}}(\ell^{\infty}(X))}\|f\|_{\mathcal{H}^{\infty}(\varphi)},$
as was to be shown.
Finally, boundedness of $C_{\psi}:\mathcal{H}^{2}(\varphi)\to\ell^{2}(X)$ is equivalent to $\psi$ being a Bessel sequence, since $\mathcal{H}^{2}(\varphi)=\mathcal{H}$. ∎
Lemma 19.
Let $\varphi=(\varphi_{k})_{k\in X}$ be a frame in $\mathcal{H}$ and $\psi=(\psi_{k})_{k\in X}$ be a family in ${\mathcal{H}}$ that satisfies $\psi\sim_{\mathcal{A}}\widetilde{\varphi}$. Then $D_{\psi}:\ell^{p}(X)\to{\mathcal{H}}^{p}(\varphi)$, defined analogously as in (13) and (9), is well-defined and bounded for each $1\leq p\leq\infty$ and the associated series converges unconditionally (resp. unconditionally in the $\sigma(\mathcal{H},\mathcal{H}^{00}(\widetilde{\varphi}))$-topology in the case $p=\infty$).
Proof.
Assume that $1\leq p<\infty$. Then, by [28, Proposition 3.22], $V^{p}(\varphi)=\{f\in\mathcal{H}:C_{\widetilde{\varphi}}f\in\ell^{p}(X)\}$ is a norm-dense subspace of $\mathcal{H}^{p}(\varphi)$. For $c=(c_{l})_{l\in X}\in\ell^{00}(X)$ we have that $D_{\psi}c\in\mathcal{H}$ as it is a finite linear combination of elements from $\psi$. Since
$\|D_{\psi}c\|_{\mathcal{H}^{p}(\varphi)}=\|C_{\widetilde{\varphi}}D_{\psi}c\|_{\ell^{p}(X)}=\|G_{\widetilde{\varphi},\psi}c\|_{\ell^{p}(X)}\leq\|G_{\widetilde{\varphi},\psi}\|_{{\mathcal{B}}(\ell^{p}(X))}\|c\|_{\ell^{p}(X)},$
we see that $D_{\psi}$ maps $\ell^{00}(X)$ boundedly into $V^{p}(\varphi)\subseteq\mathcal{H}^{p}(\varphi)$. Hence, the claim follows via density.
Now consider the case $p=\infty$. Pick some arbitrary $c=(c_{l})_{l\in X}\in\ell^{\infty}(X)$ and assume that $F_{1}\subseteq F_{2}\subseteq\dots$ is a nested sequence of finite subsets of $X$ such that $\bigcup_{n=1}^{\infty}F_{n}=X$. Then $f_{n}:=\sum_{l\in F_{n}}c_{l}\psi_{l}$ is contained in $\mathcal{H}$ for each $n\in\mathbb{N}$ and
(11) $\lim_{n\rightarrow\infty}\langle f_{n},\widetilde{\varphi}_{k}\rangle=\lim_{n\rightarrow\infty}\sum_{l\in F_{n}}\langle\psi_{l},\widetilde{\varphi}_{k}\rangle c_{l}=\sum_{l\in X}[G_{\widetilde{\varphi},\psi}]_{k,l}c_{l}=[G_{\widetilde{\varphi},\psi}c]_{k}\in\mathbb{C}$
for each $k\in X$, since $G_{\widetilde{\varphi},\psi}\in{\mathcal{A}}\subseteq{\mathcal{B}}(\ell^{\infty}(X))$ by assumption. Furthermore,
$|\langle f_{n},\widetilde{\varphi}_{k}\rangle|$ $\leq\sum_{l\in F_{n}}\big{|}[G_{\widetilde{\varphi},\psi}]_{k,l}c_{l}\big{|}$
$\leq\sup_{k\in X}\sum_{l\in X}\big{|}[G_{\widetilde{\varphi},\psi}]_{k,l}\big{|}\,|c_{l}|$
$\leq\Big{\|}\big{|}G_{\widetilde{\varphi},\psi}\big{|}\Big{\|}_{{\mathcal{B}}(\ell^{\infty}(X))}\|c\|_{\ell^{\infty}(X)}$
for all $k\in X$ and all $n\in\mathbb{N}$, where we used that $\big{|}G_{\widetilde{\varphi},\psi}\big{|}\in{\mathcal{A}}\subseteq{\mathcal{B}}(\ell^{\infty}(X))$ due to solidity of ${\mathcal{A}}$. This means that $D_{\psi}c=\big{[}(f_{n})_{n\in{\mathbb{N}}}\big{]}_{\sim_{\varphi}}\in\mathcal{H}^{\infty}(\varphi)$. The boundedness of $D_{\psi}$ follows immediately via (11), as we have
$\|D_{\psi}c\|_{\mathcal{H}^{\infty}(\varphi)}=\sup_{k\in X}\big{|}[G_{\widetilde{\varphi},\psi}c]_{k}\big{|}\leq\|G_{\widetilde{\varphi},\psi}\|_{{\mathcal{B}}(\ell^{\infty}(X))}\|c\|_{\ell^{\infty}(X)}.$
To show the unconditional convergence of $D_{\psi}c$ with respect to the $\sigma(\mathcal{H},\mathcal{H}^{00}(\widetilde{\varphi}))$-topology, it suffices to show that $D_{\psi}c$ is independent of the choice of $\{F_{n}\}_{n=1}^{\infty}$ [28, Sec. 3.2.4]. So, let $J_{1}\subseteq J_{2}\subseteq\dots$ be another nested sequence of finite subsets of $X$ with $\bigcup_{n=1}^{\infty}J_{n}=X$. We will show that $\{\sum_{l\in J_{n}}c_{l}\psi_{l}\}_{n=1}^{\infty}\sim_{\varphi}\{\sum_{l\in F_{n}}c_{l}\psi_{l}\}_{n=1}^{\infty}$. Setting $H_{n}=F_{n}\cup J_{n}$ ($n\in\mathbb{N}$) gives a nested sequence of finite subsets whose union is $X$ as well. Then, for each $k\in X$,
$\lim_{n\rightarrow\infty}\left|\left\langle\sum_{l\in H_{n}}c_{l}\psi_{l}-\sum_{l\in F_{n}}c_{l}\psi_{l},\widetilde{\varphi}_{k}\right\rangle\right|$ $=\lim_{n\rightarrow\infty}\left|\left\langle\sum_{l\in H_{n}\setminus F_{n}}c_{l}\psi_{l},\widetilde{\varphi}_{k}\right\rangle\right|$
$\leq\|c\|_{\ell^{\infty}_{\omega}(X)}\lim_{n\rightarrow\infty}\sum_{l\in X\setminus F_{n}}|\langle\psi_{l},\widetilde{\varphi}_{k}\rangle|=0,$
since
$\sup_{k\in X}\sum_{l\in X}|\langle\psi_{l},\widetilde{\varphi}_{k}\rangle|=\|G_{\widetilde{\varphi},\psi}\|_{{\mathcal{B}}(\ell^{\infty}(X))}<\infty.$
Clearly, it also holds
$\sup_{n\in\mathbb{N}}\left\|C_{\widetilde{\varphi}}\left(\sum_{l\in H_{n}}c_{l}\psi_{l}-\sum_{l\in F_{n}}c_{l}\psi_{l}\right)\right\|<\infty.$
Thus we have shown that $\{\sum_{l\in H_{n}}c_{l}\psi_{l}\}_{n=1}^{\infty}\sim_{\varphi}\{\sum_{l\in F_{n}}c_{l}\psi_{l}\}_{n=1}^{\infty}$. The same argument holds if we replace $F_{n}$ by $J_{n}$. Thus, $\{\sum_{l\in J_{n}}c_{l}\psi_{l}\}_{n=1}^{\infty}\sim_{\varphi}\{\sum_{l\in F_{n}}c_{l}\psi_{l}\}_{n=1}^{\infty}$, since $\sim_{\widetilde{\varphi}}$ is an equivalence relation. ∎
When we consider the operators $C_{\psi}:{\mathcal{H}}^{p}(\varphi)\to\ell^{p}(X)$ and $D_{\psi}:\ell^{p}(X)\to{\mathcal{H}}^{p}(\varphi)$, $1\leq p\leq\infty$, we write $\text{Ran}_{p}(C_{\psi}),\text{Ran}_{p}(D_{\psi})$ and $\text{Ker}_{p}(C_{\psi}),\text{Ker}_{p}(D_{\psi})$ for their respective ranges and nullspaces.
The next two lemmata are mild generalizations of parts of [5, Lemma 7].
Lemma 20.
Let $\varphi=(\varphi_{k})_{k\in X}$ be a frame in $\mathcal{H}$ satisfying $\varphi\sim_{{\mathcal{A}}}\varphi$ and $\psi=(\psi_{k})_{k\in X}$ be a family in ${\mathcal{H}}$ that satisfies $\psi\sim_{\mathcal{A}}\varphi$. Let $1\leq p<\infty$ be arbitrary and $\frac{1}{p}+\frac{1}{q}=1$. Then the adjoint $D_{\psi}^{\prime}:\mathcal{H}^{q}(\varphi)\to\ell^{q}(X)$ of $D_{\psi}:\ell^{p}(X)\to{\mathcal{H}}^{p}(\varphi)$ is given by $C_{\psi}$.
Proof.
We first observe that for each $1\leq p<\infty$, we have $\psi_{l}\in\mathcal{H}^{p}(\varphi)$ for each $l\in X$ due to
$\|\psi_{l}\|_{\mathcal{H}^{p}(\varphi)}=\|C_{\widetilde{\varphi}}\psi_{l}\|_{\ell^{p}(X)}\leq\|C_{\widetilde{\varphi}}\psi_{l}\|_{\ell^{1}(X)}\leq\sup_{l\in X}\sum_{k\in X}|\langle\psi_{l},\widetilde{\varphi}_{k}\rangle|=\|G_{\widetilde{\varphi},\psi}\|_{{\mathcal{B}}(\ell^{1}(X))}.$
We consider the case $(p,q)=(1,\infty)$. Let $f=\big{[}(f_{n})_{n\in{\mathbb{N}}}\big{]}_{\sim_{\varphi}}\in{\mathcal{H}}^{\infty}(\varphi)$ and $c=(c_{k})_{k\in X}\in\ell^{1}(X)$ be arbitrary. Then
$\langle c,D_{\psi}^{\prime}f\rangle_{\ell^{1}\times\ell^{\infty}}=\langle D_{\psi}c,f\rangle_{\mathcal{H}^{1}\times\mathcal{H}^{\infty}}=\sum_{k\in X}c_{k}\langle\psi_{k},f\rangle_{\mathcal{H}^{1}\times\mathcal{H}^{\infty}}.$
According to Theorem 15, we have that
$\langle\psi_{k},f\rangle_{\mathcal{H}^{1}\times\mathcal{H}^{\infty}}$ $=\sum_{l\in X}\langle\psi_{k},\widetilde{\varphi}_{l}\rangle\lim_{n\to\infty}\langle\varphi_{l},f_{n}\rangle$
$=\lim_{n\to\infty}\sum_{l\in X}\langle\psi_{k},\widetilde{\varphi}_{l}\rangle\langle\varphi_{l},f_{n}\rangle$
$=\lim_{n\to\infty}\langle\psi_{k},f_{n}\rangle=\overline{[C_{\psi}f]_{k}},$
where we applied the dominated convergence theorem (justified analogously as in (10) upon noticing that $\|C_{\varphi}\cdot\|_{\ell^{\infty}(X)}\asymp\|C_{\widetilde{\varphi}}\cdot\|_{\ell^{\infty}(X)}$ (c.f. Proposition 12)) in the second step and frame reconstruction in the third step. On the other hand,
$\langle c,C_{\psi}f\rangle_{\ell^{1}\times\ell^{\infty}}=\sum_{k\in X}c_{k}\overline{[C_{\psi}f]_{k}},$
which yields the claim. For the other values of $p$ and $q$ we proceed analogously by showing the identity for $f\in{\mathcal{H}}^{00}(\varphi)$ and then argue via density. ∎
Lemma 21.
Let $\varphi=(\varphi_{k})_{k\in X}$ be a frame in $\mathcal{H}$ satisfying $\varphi\sim_{{\mathcal{A}}}\varphi$ and $\psi=(\psi_{k})_{k\in X}$ be a family in ${\mathcal{H}}$ that satisfies $\psi\sim_{\mathcal{A}}\varphi$. Let $1\leq p<\infty$ be arbitrary and $\frac{1}{p}+\frac{1}{q}=1$. Then the Banach space adjoint $C_{\psi}^{\prime}:\ell^{q}(X)\to{\mathcal{H}}^{q}(\varphi)$ of $C_{\psi}:\mathcal{H}^{p}(\varphi)\to\ell^{p}(X)$ is given by $D_{\psi}$.
Proof.
We first consider the case $(p,q)=(1,\infty)$. Let $f\in{\mathcal{H}}^{00}(\varphi)\subseteq{\mathcal{H}}^{1}(\varphi)$ and $c=(c_{l})_{l\in X}\in\ell^{\infty}$ be arbitrary. According to Lemma 19, $D_{\psi}c=\big{[}(f_{n})_{n\in{\mathbb{N}}}\big{]}_{\sim_{\varphi}}\in{\mathcal{H}}^{\infty}(\varphi)$, where $f_{n}=\sum_{l\in F_{n}}c_{l}\psi_{l}$ and $F_{1}\subseteq F_{2}\subseteq\dots$ is a nested sequence of finite subsets of $X$ such that $\bigcup_{n=1}^{\infty}F_{n}=X$. Then we see that
$\langle f,C_{\psi}^{\prime}c\rangle_{\mathcal{H}^{1}\times\mathcal{H}^{\infty}}$ $=\langle C_{\psi}f,c\rangle_{\ell^{1}\times\ell^{\infty}}$
$=\sum_{l\in X}\langle f,\psi_{l}\rangle\overline{c_{l}}$
$=\lim_{n\to\infty}\left\langle f,\sum_{l\in F_{n}}c_{l}\psi_{l}\right\rangle$
$=\lim_{n\to\infty}\left\langle\sum_{k\in X}\langle f,\widetilde{\varphi}_{k}\rangle\varphi_{k},\sum_{l\in F_{n}}c_{l}\psi_{l}\right\rangle$
$=\sum_{k\in X}\langle f,\widetilde{\varphi}_{k}\rangle\lim_{n\to\infty}\left\langle\varphi_{k},\sum_{l\in F_{n}}c_{l}\psi_{l}\right\rangle$
$=\langle f,D_{\psi}c\rangle_{\mathcal{H}^{1}\times\mathcal{H}^{\infty}},$
where we used frame reconstruction in the fourth step and the dominated convergence theorem (justified analogously as in (10) upon noticing that $\|C_{\varphi}\cdot\|_{\ell^{\infty}(X)}\asymp\|C_{\widetilde{\varphi}}\cdot\|_{\ell^{\infty}(X)}$ (c.f. Proposition 12)) in the fifth step. Via density we conclude that $\langle f,C_{\psi}^{\prime}c\rangle_{\mathcal{H}^{1}\times\mathcal{H}^{\infty}}=\langle f,D_{\psi}c\rangle_{\mathcal{H}^{1}\times\mathcal{H}^{\infty}}$ for all $f\in\mathcal{H}^{1}(\varphi)$ and the claim is proven. For other values of $p$ and $q$ the proof is similar. ∎
As a consequence we immediately obtain the following.
Corollary 22.
Let $\varphi=(\varphi_{k})_{k\in X}$ be a frame in $\mathcal{H}$ satisfying $\varphi\sim_{{\mathcal{A}}}\varphi$ and $\psi=(\psi_{k})_{k\in X}$ be a family in ${\mathcal{H}}$ that satisfies $\psi\sim_{\mathcal{A}}\varphi$. For $1\leq p<\infty$ and $\frac{1}{p}+\frac{1}{q}=1$ the following hold.
  • (i)
    $S_{\psi}\in{\mathcal{B}}({\mathcal{H}}^{p}(\varphi))$ and $S_{\psi}^{\prime}=S_{\psi}\in{\mathcal{B}}({\mathcal{H}}^{q}(\varphi))$.
  • (ii)
    $G_{\psi}\in{\mathcal{B}}(\ell^{p}(X))$ and $G_{\psi}^{\prime}=G_{\psi}\in{\mathcal{B}}(\ell^{q}(X))$.
We conclude this subsection with the following simple consequence of Remark 3.
Lemma 23.
Let $\varphi$ be an intrinsically ${\mathcal{A}}$-localized Riesz basis in $\mathcal{H}$. Then $G_{\varphi}^{\alpha}\in{\mathcal{A}}$ for every $\alpha\in\mathbb{R}$.
Proof.
Since ${\mathcal{A}}$ is inverse-closed in ${\mathcal{B}}(\ell^{2}(X))$, the spectral identity $\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)=\sigma_{\mathcal{A}}(A)$ holds true for all $A\in{\mathcal{A}}$, and the holomorphic functional calculi associated with ${\mathcal{A}}$ and ${\mathcal{B}}(\ell^{2}(X))$ coincide (see Remark 3). In particular, if $A\leq B$ denote the Riesz basis constants of $\varphi$, then $\sigma_{{\mathcal{A}}}(G_{\varphi})=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(G_{\varphi})\subseteq[A,B]$. Thus, there exists a closed path $\Gamma$ contained in the right half plane $\text{Re}(z)>0$ with $\sigma_{{\mathcal{B}}(\ell^{2}(X))}(G_{\varphi})=\sigma_{{\mathcal{A}}}(G_{\varphi})\subset\text{int}(\Gamma)$. Since $f(z)=z^{\alpha}$ is holomorphic on $\text{Re}(z)>0$, the holomorphic functional calculus yields $G_{\varphi}^{\alpha}\in{\mathcal{A}}$. ∎
Corollary 24.
Let $\varphi$ be an intrinsically ${\mathcal{A}}$-localized Riesz basis in $\mathcal{H}$. Then both $G_{S_{\varphi}^{-1/4}\varphi}$ and $G_{S_{\varphi}^{-3/4}\varphi}$ are contained in ${\mathcal{A}}$.
Proof.
Since $\varphi$ is a Riesz basis, $(S_{\varphi}^{-1/2}\varphi_{k})_{k\in X}$ is an orthonormal basis. By a simple calculation we see that $\big{(}G_{S_{\varphi}^{-1/4}\varphi}\big{)}^{2}=G_{\varphi}$. Thus, by Lemma 23, $G_{S_{\varphi}^{-1/4}\varphi}=G_{\varphi}^{1/2}\in{\mathcal{A}}$. Similarly, we have $(G_{S_{\varphi}^{-3/4}\varphi})^{2}=G_{\widetilde{\varphi}}\in{\mathcal{A}}$ due to Theorem 6, which, by Lemma 23, implies $G_{S_{\varphi}^{-3/4}\varphi}\in\mathcal{A}$. ∎
3.2. Localization of the R-dual sequence
One main ingredient in Theorem 1 is the family $\omega$, which, in the language of [33], is a so-called R-dual of type III of $\psi$.
Let $(\psi_{k})_{k\in X}$ be a Bessel sequence in $\mathcal{H}$ and let $\varphi=(\varphi_{k})_{k\in X}$ be a Riesz basis in $\mathcal{H}$. Then $(S_{\varphi}^{-1/2}\varphi_{k})_{k\in X}$ is an orthonormal basis in $\mathcal{H}$ [12]. Consequently,
$\omega_{k}=\sum_{l\in X}\langle\psi_{l},\varphi_{k}\rangle S_{\varphi}^{-{1}/{2}}\varphi_{l}$
converges in $\mathcal{H}$ for each $k\in X$, defining a family $\omega=(\omega_{k})_{k\in X}$ of vectors in $\mathcal{H}$. One important observation for proving proving Theorem 1 is, that $\omega\sim_{{\mathcal{A}}}\varphi$ whenever $\psi\sim_{{\mathcal{A}}}\varphi$, see below.
Lemma 25.
Let ${\mathcal{A}}$ be a spectral algebra and $\varphi=(\varphi_{k})_{k\in X}$ be a Riesz basis in $\mathcal{H}$ that satisfies $\varphi\sim_{{\mathcal{A}}}\varphi$. Let $\psi=(\psi_{k})_{k\in X}$ be a sequence in $\mathcal{H}$ satisfying $\psi\sim_{{\mathcal{A}}}\varphi$ and let $\omega=(\omega_{k})_{k\in X}$ be given by
$\omega_{k}=\sum_{l\in X}\langle\psi_{l},\varphi_{k}\rangle S_{\varphi}^{-{1}/{2}}\varphi_{l}\qquad(k\in X).$
Then $\omega\sim_{{\mathcal{A}}}\varphi$, $\omega\sim_{{\mathcal{A}}}\widetilde{\varphi}$ and $\omega\sim_{{\mathcal{A}}}\omega$. In particular, $\omega$ is a Bessel sequence in $\mathcal{H}$.
Proof.
First, note that $\psi$ is a Bessel sequence, since $G_{\psi}\in{\mathcal{A}}\subset{\mathcal{B}}(\ell^{2}(X))$ according to Lemma 7, whence $\omega$ is a well-defined sequence in $\mathcal{H}$. Next, recall from Corollary 24 that $G_{S_{\varphi}^{-1/4}\varphi}\in{\mathcal{A}}$. Since
$\langle\varphi_{n},\omega_{k}\rangle=\sum_{l\in X}\langle\varphi_{k},\psi_{l}\rangle\left\langle S_{\varphi}^{-1/4}\varphi_{n},S_{\varphi}^{-1/4}\varphi_{l}\right\rangle=\sum_{l\in X}[G_{\psi,\varphi}^{*}]_{k,l}\left[G_{S_{\varphi}^{-1/4}\varphi}\right]_{l,n}$
for all $k,n\in X$, we see that $G_{\omega,\varphi}=G_{\psi,\varphi}^{*}G_{S_{\varphi}^{-1/4}\varphi}\in{\mathcal{A}}$.
Similarly, since $G_{S_{\varphi}^{-3/4}\varphi}\in\mathcal{A}$ and
$\langle\widetilde{\varphi}_{n},\omega_{k}\rangle=\sum_{l\in X}\langle\varphi_{k},\psi_{l}\rangle\left\langle S_{\varphi}^{-3/4}\varphi_{n},S_{\varphi}^{-3/4}\varphi_{l}\right\rangle=\sum_{l\in X}[G_{\psi,\varphi}^{*}]_{k,l}\left[G_{S_{\varphi}^{-3/4}\varphi}\right]_{l,n}$
for all $k,n\in X$, we obtain $G_{\omega,\widetilde{\varphi}}=G_{\psi,\varphi}^{*}G_{S_{\varphi}^{-3/4}\varphi}\in{\mathcal{A}}$.
Finally, $G_{\omega}=G_{\omega,\widetilde{\varphi}}G_{{\varphi},\omega}\in{\mathcal{A}}$, hence $G_{\omega}\in{\mathcal{B}}(\ell^{2}(X))$, which implies that $\omega$ is a Bessel sequence in $\mathcal{H}$. ∎
As a consequence, all results appearing in the previous subsection apply to $\omega$, which will be crucial for proving Theorem 1.
3.3. Complex Interpolation
Computing the complex interpolation spaces for a couple of closed subspaces $(X_{0},X_{1})\subset(L^{p_{0}}(\Omega),L^{p_{1}}(\Omega))$ is nontrivial and solutions for certain special cases rely, for example, on the complemented subspace property [35, Theorem 1.17.1]. For our purposes it will however be sufficient to observe that the complex interpolation spaces $(X_{0},X_{1})_{\theta}$, $\theta\in(0,1)$ are continuously embedded in $L^{p_{\theta}}(\Omega)$, $\frac{1}{p_{\theta}}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}$. The proof is essentially a repetition of one direction of the proof of the standard complex interpolation result for $L^{p}(\Omega)$ spaces. For the convenience of the reader we repeat it here and refer to [9] for a detailed exposition on interpolation spaces.
Lemma 26.
Let $(\Omega,\mu)$ be a measure space with nonnegative measure $\mu$, $L^{p}(\Omega)=L^{p}(\Omega,\mu)$, and $X_{0}\subset L^{p_{0}}(\Omega)$ and $X_{1}\subset L^{p_{1}}(\Omega)$, $1\leq p_{0},p_{1}\leq\infty$, be closed subspaces that form a compatible couple $(X_{0},X_{1})$ with $X_{0}\subset X_{1}$. Moreover, let $\theta\in(0,1)$ and $\frac{1}{p_{\theta}}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}$. Then $L^{p_{\theta}}(\Omega)\cap X_{0}\hookrightarrow(X_{0},X_{1})_{\theta}$, that is
$\|f\|_{\theta}\leq\|f\|_{L^{p_{\theta}}(\Omega)},\qquad f\in X_{0}.$
Proof.
First, note that $X_{0}\cap X_{1}=X_{0}$ and $X_{0}+X_{1}=X_{1}$. Define $\mathcal{F}(X_{0},X_{1})$ as the space of $X_{1}$-valued functions that are holomorphic on the strip $S=\{z\in\mathbb{C}:0<\text{Re}(z)<1\}$, continuous on its closure and satisfy $\{F(it):\ t\in\mathbb{R}\}\subset X_{0},$ as well as $\{F(1+it):\ t\in\mathbb{R}\}\subset X_{1}$. When equipped with the norm
$\|F\|_{\mathcal{F}(X_{0},X_{1})}:=\max\Big{\{}\sup_{t\in\mathbb{R}}\|F(it)\|_{X_{0}},\sup_{t\in\mathbb{R}}\|F(1+{it})\|_{X_{1}}\Big{\}},$
$\mathcal{F}(X_{0},X_{1})$ is a Banach space. The complex interpolation space $(X_{0},X_{1})_{\theta}$, $\theta\in(0,1)$, is then given by
$(X_{0},X_{1})_{\theta}=\big{\{}f\in X_{1}:\ f=F(\theta),\ F\in\mathcal{F}(X_{0},X_{1})\big{\}}$
equipped with the norm
$\|f\|_{\theta}=\inf\big{\{}\|F\|_{\mathcal{F}(X_{0},X_{1})}:f=F(\theta),\ F\in\mathcal{F}(X_{0},X_{1})\big{\}}.$
Let $f\in X_{0}$. Since $X_{0}\subset L^{p_{0}}(\Omega)\cap L^{p_{1}}(\Omega)$ we may without loss of generality assume that $\|f\|_{L^{p_{\theta}}(\Omega)}=1$ and choose $F^{\ast}(z)=|f|^{p_{\theta}/p(z)-1}f$, where $\frac{1}{p(z)}=\frac{1-z}{p_{0}}+\frac{z}{p_{1}}$. Clearly $F^{\ast}\in\mathcal{F}(X_{0},X_{1})$, $F^{\ast}(\theta)=f$, and
$\|f\|_{\theta}$ $\leq\|F^{\ast}\|_{\mathcal{F}(X_{0},X_{1})}$
$=\max\Big{\{}\sup_{t\in\mathbb{R}}\big{\|}|f|^{p_{\theta}(\frac{1-it}{p_{0}}+\frac{it}{p_{1}})-1}f\big{\|}_{L^{p_{0}}(\Omega)},\sup_{t\in\mathbb{R}}\big{\|}|f|^{p_{\theta}(-\frac{it}{p_{0}}+\frac{1+it}{p_{1}})-1}f\big{\|}_{L^{p_{1}}(\Omega)}\Big{\}}$
$=\max\Big{\{}\|f\|_{L^{p_{\theta}}(\Omega)}^{1/p_{0}},\|f\|_{L^{p_{\theta}}(\Omega)}^{1/p_{1}}\Big{\}}=1=\|f\|_{L^{p_{\theta}}(\Omega)}.$
4. Proof of Theorem 1
12356741089
Figure 1. Graph showing which implications of the equivalent statements in Theorem 1 are shown.
Proof.
We show the equivalences of Theorem 1 via a chain of implications that is depicted in Figure 1. Before we start, observe that by Lemma 25 as well as Lemma 18 and Lemma 19, all operators occurring in the statements (2)--(9) are bounded on their respective domains.
(1$\Rightarrow$(2): The assumptions imply that $\psi$ is an intrinsically ${\mathcal{A}}$-localized frame, since $G_{\psi}=G_{\psi,\varphi}G_{\widetilde{\varphi}}G_{\varphi,\psi}\in{\mathcal{A}}$. Consequently, by Theorem 14, $S_{\psi}$ is bounded and invertible on $\mathcal{H}^{1}(\psi)$. By Remark 9 it holds that $\mathcal{H}^{1}(\psi)=\mathcal{H}^{1}(\varphi)$ with equivalent norms. Hence (2) follows.
(2$\Leftrightarrow$(3): This follows from duality via Corollary 22 (i).
(2$\Rightarrow$(5): Since $S_{\psi}=D_{\psi}C_{\psi}:{\mathcal{H}}^{1}(\varphi)\to{\mathcal{H}}^{1}(\varphi)$ is bijective, $D_{\psi}:\ell^{1}(X)\to{\mathcal{H}}^{1}(\varphi)$ has to be surjective and $C_{\psi}:{\mathcal{H}}^{1}(\varphi)\to\ell^{1}(X)$ injective.
To prove the remaining part of the statement, let us assume to the contrary that $\text{Ran}_{1}(C_{\psi})$ is not closed in $\ell^{1}(X)$. Then $D_{\psi}:\overline{\text{Ran}_{1}(C_{\psi})}\to{\mathcal{H}}^{1}(\varphi)$ cannot be injective since $D_{\psi}:\text{Ran}_{1}(C_{\psi})\to{\mathcal{H}}^{1}(\varphi)$ is necessarily bijective. Therefore, we may assume that there exists some $c\in\overline{\text{Ran}_{1}(C_{\psi})}\setminus\{0\}$ such that $D_{\psi}c=0$ and $\|c\|_{\ell^{1}(X)}=1$. Now let $(c^{(n)})_{n\in{\mathbb{N}}}\subset{\text{Ran}_{1}(C_{\psi})}$ be a sequence with $\|c^{(n)}\|_{\ell^{1}(X)}=1$ for each $n\in\mathbb{N}$, which converges in $\ell^{1}(X)$ to $c$. By continuity of $D_{\psi}$ (Lemma 19) we obtain that $D_{\psi}c^{(n)}\to 0$ in $\mathcal{H}^{1}(\varphi)$ as $n\to\infty$. If $(f_{n})_{n\in{\mathbb{N}}}\subset\ell^{1}(X)$ is the sequence with $C_{\psi}f_{n}=c^{(n)}$ (for each $n\in\mathbb{N}$), then we see that
$\lim_{n\to\infty}\|f_{n}\|_{{\mathcal{H}}^{1}(\varphi)}\lesssim\lim_{n\to\infty}\|S_{\psi}f_{n}\|_{{\mathcal{H}}^{1}(\varphi)}=\lim_{n\to\infty}\|D_{\psi}c^{(n)}\|_{{\mathcal{H}}^{1}(\varphi)}=0.$
Since $G_{\psi,\varphi}\in{\mathcal{A}}\subset{\mathcal{B}}(\ell^{1}(X))$, this finally implies that
$\lim_{n\to\infty}\|c^{(n)}\|_{\ell^{1}(X)}=\lim_{n\to\infty}\|G_{\psi,\varphi}C_{\widetilde{\varphi}}f_{n}\|_{\ell^{1}(X)}\lesssim\lim_{n\to\infty}\|f_{n}\|_{{\mathcal{H}}^{1}(\varphi)}=0,$
a contradiction. Consequently, $\text{Ran}_{1}(C_{\psi})$ must be closed in $\ell^{1}(X)$.
(4$\Leftrightarrow$(5): By duality, Lemma 20 and the closed range theorem, $C_{\psi}:\mathcal{H}^{\infty}(\varphi)\to\ell^{\infty}(X)$ is injective and has closed range if and only if $D_{\psi}:\ell^{1}(X)\to\mathcal{H}^{1}(\varphi)$ is surjective. By analogous reasoning via Lemma 21, $\text{Ran}_{1}(C_{\psi})$ is closed if and only $\text{Ran}_{\infty}(D_{\psi})$ is closed.
(5$\Rightarrow$(6): First, we show that $D_{\omega}:\ell^{\infty}\to\mathcal{H}^{\infty}(\varphi)$ is injective. By Theorem 14, $S_{\varphi}\in{\mathcal{B}}(\mathcal{H}^{\infty}(\varphi))$ is bounded and invertible, i.e., 00 is contained in the resolvent set of $S_{\varphi}$ with respect to ${\mathcal{B}}(\mathcal{H}^{\infty}(\varphi))$. Since $z\mapsto z^{1/2}$ is holomorphic on any open subset $U\subset\mathbb{C}$ that does not contain the origin, we may deduce from the holomorphic functional calculus that $S_{\varphi}^{{1}/{2}}\in{\mathcal{B}}({\mathcal{H}}^{\infty}(\varphi))$ is bounded and invertible as well. Now let $c\in\ell^{\infty}(X)$ be such that $0=D_{\omega}c$ in $\mathcal{H}^{\infty}(\varphi)$. By definition, this means that $\lim_{n\to\infty}\langle[D_{\omega}c]_{n},\widetilde{\varphi}_{k}\rangle=0$ for each $k\in X$. Letting $F_{1}\subseteq F_{2}\subseteq\dots$ be a nested sequence of finite subsets of $X$ with $\bigcup_{n=1}^{\infty}F_{n}=X$, and recalling from Corollary 24 that $G_{S_{\varphi}^{-3/4}\varphi}\in{\mathcal{A}}\subseteq{\mathcal{B}}(\ell^{\infty}(X))$, we see that
00 0 $=\lim_{n\to\infty}\left\langle\sum_{l\in F_{n}}c_{l}\omega_{l},\widetilde{\varphi}_{k}\right\rangle$
$=\sum_{l\in X}c_{l}\left\langle\sum_{m\in X}\langle\psi_{m},\varphi_{l}\rangle S_{\varphi}^{-1/2}\varphi_{m},\widetilde{\varphi}_{k}\right\rangle$
$=\sum_{m\in X}\langle S_{\varphi}^{-1/2}\varphi_{m},\widetilde{\varphi}_{k}\rangle\sum_{l\in X}c_{l}\langle\psi_{m},\varphi_{l}\rangle$
(12) $=\sum_{m\in X}\big{[}G_{S_{\varphi}^{-3/4}\varphi}\big{]}_{k,m}\left[\,\overline{C_{\psi}}D_{\varphi}\overline{c}\right]_{m}\qquad(\text{for each }k\in X),$
where $\overline{C_{\psi}}:{\mathcal{H}}^{\infty}(\varphi)\to\ell^{\infty}(X)$ is defined by $\overline{C_{\psi}}f=\big{(}\,\overline{[C_{\psi}f]_{k}}\,\big{)}_{k\in X}$. Note that we are allowed to interchange the order of summation since
$\sup_{k\in X}\sum_{m\in X}\big{|}G_{S_{\varphi}^{-3/4}\varphi}\big{|}_{k,m}\cdot\sup_{s\in X}\sum_{l\in X}\big{|}G_{\varphi,\psi}\big{|}_{s,l}<\infty,$
which follows from $G_{S_{\varphi}^{-3/4}\varphi},G_{\varphi,\psi}\in{\mathcal{A}}\subseteq{\mathcal{B}}(\ell^{\infty}(X))$. From (12) we know that $G_{S_{\varphi}^{-3/4}\varphi}\overline{C_{\psi}}D_{\varphi}\overline{c}=0$ in $\ell^{\infty}(X)$. By Lemma 16, $G_{\widetilde{\varphi}}\in{\mathcal{A}}\subset{\mathcal{B}}(\ell^{\infty}(X))$ is invertible. Arguing as in the proof of Lemma 25, we apply the holomorphic functional calculus to show that $G_{S_{\varphi}^{-3/4}\varphi}=(G_{\widetilde{\varphi}})^{1/2}\in{\mathcal{A}}\subset{\mathcal{B}}(\ell^{\infty}(X))$ is invertible as well. Consequently, $\overline{C_{\psi}}D_{\varphi}\overline{c}=0$. By (4$\Leftrightarrow$(5), $C_{\psi}:{\mathcal{H}}^{\infty}(\varphi)\to\ell^{\infty}(X)$ is injective, which is equivalent to $\overline{C_{\psi}}:{\mathcal{H}}^{\infty}(\varphi)\to\ell^{\infty}(X)$ being injective. Consequently, $D_{\varphi}\overline{c}=0$. Finally, an application of Corollary 17 yields $c=0$.
Next, we show that $\text{Ran}_{\infty}(D_{\omega})$ is closed. Let $(D_{\omega}c^{(m)})_{m\in{\mathbb{N}}}\subset\mathcal{H}^{\infty}(\varphi)$ be a sequence in $\text{Ran}_{\infty}(D_{\omega})$ which converges in $\mathcal{H}^{\infty}(\varphi)$ to some $f=\big{[}(f_{n})_{n\in{\mathbb{N}}}\big{]}_{\sim_{\varphi}}\in\mathcal{H}^{\infty}(\varphi)$. In particular, for each $m\in\mathbb{N}$, we have by definition that $D_{\omega}c^{(m)}=\big{[}\big{(}\sum_{l\in F_{n}^{(m)}}c_{l}^{(m)}\omega_{l}\big{)}_{n\in{\mathbb{N}}}\big{]}_{\sim_{\varphi}}$, where $F_{1}^{(m)}\subseteq F_{2}^{(m)}\subseteq\dots$ is a nested sequence of finite subsets of $X$ so that $X=\bigcup_{n\in F_{n}^{(m)}}$. By Lemma 19, we may assume without loss of generality that for each $n\in\mathbb{N}$, the sets $F_{n}^{(m)}$ ($m\in\mathbb{N}$) are equal. For simplicity, we write $F_{n}=F_{n}^{(m)}$. Thus we have
00 0 $=\lim_{m\to\infty}\left\|D_{\omega}c^{(m)}-f\right\|_{\mathcal{H}^{\infty}(\varphi)}$
$=\lim_{m\to\infty}\sup_{k\in X}\lim_{n\to\infty}\left|\left\langle\sum_{l\in F_{n}}c_{l}^{(m)}\omega_{l}-f_{n},\widetilde{\varphi}_{k}\right\rangle\right|$
$=\lim_{m\to\infty}\sup_{k\in X}\lim_{n\to\infty}\left|\left\langle\sum_{l\in F_{n}}c_{l}^{(m)}\sum_{i\in X}\langle\psi_{i},\varphi_{l}\rangle S_{\varphi}^{-1/2}\varphi_{i}-f_{n},\widetilde{\varphi}_{k}\right\rangle\right|$
$=\lim_{m\to\infty}\sup_{k\in X}\lim_{n\to\infty}\left|\sum_{i\in X}\sum_{l\in F_{n}}\langle\psi_{i},\varphi_{l}\rangle\langle S_{\varphi}^{-1/2}\varphi_{i},\widetilde{\varphi}_{k}\rangle c_{l}^{(m)}-[C_{\widetilde{\varphi}}f]_{k}\right|$
$=\lim_{m\to\infty}\sup_{k\in X}\left|\sum_{i\in X}\langle S_{\varphi}^{-1/2}\varphi_{i},\widetilde{\varphi}_{k}\rangle\lim_{n\to\infty}\sum_{l\in F_{n}}\langle\psi_{i},\varphi_{l}\rangle c_{l}^{(m)}-[C_{\widetilde{\varphi}}f]_{k}\right|$
$=\lim_{m\to\infty}\sup_{k\in X}\left|\sum_{i\in X}\langle S_{\varphi}^{-1/2}\varphi_{i},\widetilde{\varphi}_{k}\rangle\lim_{n\to\infty}\left\langle\psi_{i},\Big{[}D_{\varphi}\overline{c^{(m)}}\Big{]}_{n}\right\rangle-[C_{\widetilde{\varphi}}f]_{k}\right|$
$=\lim_{m\to\infty}\sup_{k\in X}\left|\sum_{i\in X}\langle S_{\varphi}^{-1/2}\varphi_{i},\widetilde{\varphi}_{k}\rangle\left[\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}\right]_{i}-[C_{\widetilde{\varphi}}f]_{k}\right|$
$=\lim_{m\to\infty}\left\|C_{\widetilde{\varphi}}S_{\varphi}^{-1/2}D_{\varphi}\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}-C_{\widetilde{\varphi}}f\right\|_{\ell^{\infty}(X)},$
where we applied the dominated convergence theorem in the fifth step. Since $C_{\widetilde{\varphi}}S_{\varphi}^{1/2}D_{\varphi}\in{\mathcal{B}}(\ell^{\infty}(X))$, the latter implies that
$0=\lim_{m\to\infty}\left\|\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}-C_{\widetilde{\varphi}}S_{\varphi}^{1/2}f\right\|_{\ell^{\infty}(X)}.$
This means that $\big{(}\,\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}\,\big{)}_{m\in{\mathbb{N}}}$ is a sequence in $\text{Ran}_{\infty}\left(\overline{C_{\psi}}\right)$ which converges in $\ell^{\infty}(X)$ to $C_{\widetilde{\varphi}}S_{\varphi}^{1/2}f\in\ell^{\infty}(X)$. Since $\overline{C_{\psi}}$ has closed range due to $C_{\psi}$ having closed range by (4$\Leftrightarrow$(5), we must have $C_{\widetilde{\varphi}}S_{\varphi}^{1/2}f=\overline{C_{\psi}}h$ for some $h\in\mathcal{H}^{\infty}(\varphi)$. Since $D_{\varphi}:\ell^{\infty}(X)\to\mathcal{H}^{\infty}(\varphi)$ is surjective by Proposition 13, there exists some $\overline{c}\in\ell^{\infty}(X)$ with $h=D_{\varphi}\overline{c}$. Altogether, we obtain
$0=\lim_{m\to\infty}\left\|\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}-\overline{C_{\psi}}D_{\varphi}\overline{c}\right\|_{\ell^{\infty}(X)}.$
Since $C_{\widetilde{\varphi}}S_{\varphi}^{-1/2}D_{\varphi}\in{\mathcal{B}}(\ell^{\infty}(X))$, the latter yields
$0=\lim_{m\to\infty}\left\|C_{\widetilde{\varphi}}S_{\varphi}^{-1/2}\ D_{\varphi}\overline{C_{\psi}}D_{\varphi}(\overline{c^{(m)}}-\overline{c})\right\|_{\ell^{\infty}(X)}.$
By going through the calculations from above in reverse order, we arrive at
$0=\lim_{m\to\infty}\left\|D_{\omega}c^{(m)}-D_{\omega}c\right\|_{\mathcal{H}^{\infty}(\varphi)},$
which yields the claim.
Finally, we show that $\text{Ran}_{\infty}(C_{\omega})$ is closed by applying a similar argument. By Lemma 20 and the closed range theorem, it suffices to prove that $D_{\omega}:\ell^{1}(X)\to\mathcal{H}^{1}(\varphi)$ has closed range. To this end, let $(D_{\omega}c^{(m)})_{m\in{\mathbb{N}}}\subset\mathcal{H}^{1}(\varphi)$ be a sequence in $\text{Ran}_{1}(D_{\omega})$ which converges in $\mathcal{H}^{1}(\varphi)$ to some $f\in\mathcal{H}^{1}(\varphi)$. Next, employing the holomorphic functional calculus with respect to ${\mathcal{B}}(\mathcal{H}^{1}(\varphi))$, we see that $S_{\varphi}^{1/2}\in{\mathcal{B}}(\mathcal{H}^{1}(\varphi))$ is bounded and invertible. Analogously to the previous part of the proof, we calculate
00 0 $=\lim_{m\to\infty}\left\|D_{\omega}c^{(m)}-f\right\|_{\mathcal{H}^{1}(\varphi)}$
$=\lim_{m\to\infty}\sum_{k\in X}\left|\left\langle\sum_{l\in X}c_{l}^{(m)}\omega_{l},\widetilde{\varphi}_{k}\right\rangle-[C_{\widetilde{\varphi}}f]_{k}\right|$
$=\lim_{m\to\infty}\sum_{k\in X}\left|\left\langle\sum_{l\in X}c_{l}^{(m)}\sum_{i\in X}\langle\psi_{i},\varphi_{l}\rangle S_{\varphi}^{-1/2}\varphi_{i},\widetilde{\varphi}_{k}\right\rangle-[C_{\widetilde{\varphi}}f]_{k}\right|$
$=\lim_{m\to\infty}\sum_{k\in X}\left|\sum_{i\in X}\langle S_{\varphi}^{-1/2}\varphi_{i},\widetilde{\varphi}_{k}\rangle\sum_{l\in X}\langle\psi_{i},\varphi_{l}\rangle c_{l}^{(m)}-[C_{\widetilde{\varphi}}f]_{k}\right|$
$=\lim_{m\to\infty}\left\|C_{\widetilde{\varphi}}S_{\varphi}^{-1/2}D_{\varphi}\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}-C_{\widetilde{\varphi}}f\right\|_{\ell^{1}(X)}.$
Continuity of $C_{\widetilde{\varphi}}S_{\varphi}^{1/2}D_{\varphi}\in{\mathcal{B}}(\ell^{1}(X))$ implies
$0=\lim_{m\to\infty}\left\|\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}-C_{\widetilde{\varphi}}S_{\varphi}^{1/2}f\right\|_{\ell^{1}(X)}.$
This means that $\big{(}\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}\big{)}_{m\in{\mathbb{N}}}$ is a sequence in $\text{Ran}_{1}\left(\overline{C_{\psi}}\right)$ which converges in $\ell^{1}(X)$ to $C_{\widetilde{\varphi}}S_{\varphi}^{1/2}f\in\ell^{1}(X)$. Since $\text{Ran}_{1}\left(\overline{C_{\psi}}\right)$ is closed due to $\text{Ran}_{1}\left(C_{\psi}\right)$ being closed, we must have $C_{\widetilde{\varphi}}S_{\varphi}^{1/2}f=\overline{C_{\psi}}h$ for some $h\in\mathcal{H}^{1}(\varphi)$. Since $D_{\varphi}:\ell^{1}(X)\to\mathcal{H}^{1}(\varphi)$ is surjective due to Proposition 13, there exists some $\overline{c}\in\ell^{1}(X)$ with $h=D_{\varphi}\overline{c}$. Altogether, we obtain
$0=\lim_{m\to\infty}\left\|\overline{C_{\psi}}D_{\varphi}\overline{c^{(m)}}-\overline{C_{\psi}}D_{\varphi}\overline{c}\right\|_{\ell^{1}(X)}.$
Since $C_{\widetilde{\varphi}}S_{\varphi}^{-1/2}D_{\varphi}\in{\mathcal{B}}(\ell^{1}(X))$, the latter yields
$0=\lim_{m\to\infty}\left\|C_{\widetilde{\varphi}}S_{\varphi}^{-1/2}\ D_{\varphi}\overline{C_{\psi}}D_{\varphi}(\overline{c^{(m)}}-\overline{c})\right\|_{\ell^{1}(X)}.$
Finally, by going through the calculations from above in reverse order, we finally arrive at
$0=\lim_{m\to\infty}\left\|D_{\omega}c^{(m)}-D_{\omega}c\right\|_{\mathcal{H}^{1}(\varphi)},$
as was to be shown.
(6$\Leftrightarrow$(7): Lemma 25 ensures that we may apply Lemma 21 to deduce that the Banach space adjoint of $C_{\omega}:\mathcal{H}^{1}(\varphi)\to\ell^{1}(X)$ is $D_{\omega}:\ell^{\infty}(X)\to\mathcal{H}^{\infty}(\varphi)$. Therefore, $\text{Ker}_{\infty}(D_{\omega})=\{0\}$ if and only if $\text{Ran}_{1}(C_{\omega})$ is dense in $\ell^{1}(X)$. The equivalence of the conditions on the closedness of $\text{Ran}_{\infty}(C_{\omega})$ and $\text{Ran}_{1}(D_{\omega})$ as well as $\text{Ran}_{\infty}(D_{\omega})$ and $\text{Ran}_{1}(C_{\omega})$ is due to the closed range theorem.
(7$\Rightarrow$(10): Since $\text{Ran}_{1}(C_{\omega})=\ell^{1}(X)$, it follows that $\text{Ran}_{1}(C_{\omega})$ is dense in $\ell^{2}(X)$ and so is $\text{Ran}_{2}(C_{\omega})$. We thus deduce that $\text{Ker}_{1}(D_{\omega})\subset\text{Ker}_{2}(D_{\omega})=\{0\}$, that is, $D_{\omega}:\ell^{1}(X)\to\text{Ran}_{1}(D_{\omega})$ is bijective. Since $\text{Ran}_{1}(D_{\omega})$ is closed, it follows from the bounded inverse theorem that there exist constants $A_{1},B_{1}>0$ such that
(13) $A_{1}\|c\|_{\ell^{1}(X)}\leq\|D_{\omega}c\|_{{\mathcal{H}}^{1}(\varphi)}\leq B_{1}\|c\|_{\ell^{1}(X)},\qquad c\in\ell^{1}(X).$
Next, observe that by (6$\Leftrightarrow$(7), $D_{\omega}:\ell^{\infty}(X)\to{\mathcal{H}}^{\infty}(\varphi)$ is injective and has closed range. We may therefore apply the bounded inverse theorem again to obtain constants $A_{\infty},B_{\infty}>0$ such that
(14) $A_{\infty}\|c\|_{\ell^{\infty}(X)}\leq\|D_{\omega}c\|_{{\mathcal{H}}^{\infty}(\varphi)}\leq B_{\infty}\|c\|_{\ell^{\infty}(X)},\qquad c\in\ell^{\infty}(X).$
Since $C_{\omega}:{\mathcal{H}}^{1}(\varphi)\to\ell^{1}(X)$ is surjective, there exists $(\gamma_{k})_{k\in X}\subset{\mathcal{H}}^{1}(\varphi)$ such that
$\langle\gamma_{k},\omega_{l}\rangle=\delta_{k,l},\qquad k,l\in X.$
Applying (13) and (14) respectively allows us to infer that for $p\in\{1,\infty\}$
$B^{-1}_{p}\|D_{\omega}c\|_{{\mathcal{H}}^{p}(\varphi)}\leq\|c\|_{\ell^{p}(X)}=\|C_{\gamma}D_{\omega}c\|_{\ell^{p}(X)}=\|c\|_{\ell^{p}(X)}\leq A^{-1}_{p}\|D_{\omega}c\|_{{\mathcal{H}}^{p}(\varphi)}.$
In particular,
$\|C_{\gamma}f\|_{\ell^{p}(X)}\leq A^{-1}_{p}\|f\|_{{\mathcal{H}}^{p}(\varphi)},\qquad\text{for every }f\in\text{Ran}_{p}(D_{\omega}),\text{ and }p\in\{1,\infty\}.$
Note that $\big{(}\text{Ran}_{1}(D_{\omega}),\text{Ran}_{\infty}(D_{\omega})\big{)}$ forms a compatible pair of Banach spaces with $\text{Ran}_{1}(D_{\omega})\subset\text{Ran}_{\infty}(D_{\omega})$. It follows from complex interpolation that
$\|C_{\gamma}f\|_{\ell^{p_{\theta}}(X)}\leq A_{1}^{\theta-1}A_{\infty}^{-\theta}\|f\|_{\theta},\qquad\text{ for every }f\in\text{Ran}_{1}(D_{\omega}),\text{ and }0\leq\theta\leq 1.$
Since $\text{Ran}_{p}(D_{\omega})$ can be identified with a closed subspace of $\ell^{p}(X)$ via $C_{\widetilde{\varphi}}$, we may apply Lemma 26 to deduce that
$\|f\|_{\theta}=\|C_{\widetilde{\varphi}}f\|_{\theta}\leq\|C_{\widetilde{\varphi}}f\|_{\ell^{p_{\theta}}(X)}=\|f\|_{{\mathcal{H}}^{p_{\theta}}(\varphi)}.$
Specifying $\theta=1/2$ we have $p_{1/2}=2$ and applying a density argument then yields
$\|C_{\gamma}f\|_{\ell^{2}(X)}\lesssim\|f\|_{{\mathcal{H}}},\qquad f\in\text{Ran}_{2}(D_{\omega}).$
Finally,
$\|c\|_{\ell^{2}(X)}=\|C_{\gamma}D_{\omega}c\|_{\ell^{2}(X)}\lesssim\|D_{\omega}c\|_{{\mathcal{H}}}\lesssim\|c\|_{\ell^{2}(X)},\qquad c\in\ell^{2}(X),$
where the inequality on the right hand side is a consequence of Lemma 25. In other words, $\omega$ is a Riesz sequence in $\mathcal{H}$.
(10$\Leftrightarrow$(1): Note that $\big{(}S_{\varphi}^{1/2}\psi_{k}\big{)}_{k\in X}$ is a Bessel sequence because $(\psi_{k})_{k\in X}$ is a Bessel sequence by Lemma 18 and $S_{\varphi}^{1/2}\in{\mathcal{B}}(\mathcal{H})$. Since $S_{\varphi}^{1/2}$ is invertible (see, e.g., [12]), we may rewrite $\omega_{k}$ as
$\omega_{k}=\sum_{l\in X}\langle S_{\varphi}^{1/2}\psi_{l},S_{\varphi}^{-1/2}\varphi_{k}\rangle S_{\varphi}^{-1/2}\varphi_{l},$
which means that $(\omega_{k})_{k\in X}$ is an R-dual sequence of $\big{(}S_{\varphi}^{1/2}\psi_{k}\big{)}_{k\in X}$ in the sense of [11]. By [11, Theorem 2] we know that $(\omega_{k})_{k\in X}$ is a Riesz sequence if and only if $\big{(}S_{\varphi}^{1/2}\psi_{k}\big{)}_{k\in X}$ is a frame, which in turn is equivalent to $(\psi_{k})_{k\in X}$ being a frame, since $S_{\varphi}^{1/2}\in{\mathcal{B}}(\mathcal{H})$ is invertible.
(10$\Rightarrow$(9): This implication follows directly from Lemma 16, since we have that $\omega\sim_{\mathcal{A}}\omega$ by Lemma 25.
[…]
(8$\Rightarrow$(7): Since $G_{\omega}=C_{\omega}D_{\omega}:\ell^{1}(X)\to\ell^{1}(X)$ is invertible, it follows that $C_{\omega}:\mathcal{H}^{1}(\varphi)\to\ell^{1}(X)$ is surjective.
To prove the second statement, let us assume to the contrary that $\text{Ran}_{1}(D_{\omega})$ is not closed in ${\mathcal{H}}^{1}(\varphi)$. Then $C_{\omega}:\overline{\text{Ran}_{1}(D_{\omega})}\to\ell^{1}(X)$ cannot be injective. Indeed, $C_{\omega}:\text{Ran}_{1}(D_{\omega})\to\ell^{1}(X)$ is necessarily bijective as $G_{\omega}$ is invertible. We may therefore assume that there exists some $f\in\overline{\text{Ran}_{1}(D_{\omega})}\setminus\{0\}$ such that $C_{\omega}f=0$ and $\|f\|_{{\mathcal{H}}^{1}(\varphi)}=1$. Now let $\{f_{n}\}_{n=1}^{\infty}\subset{\text{Ran}_{1}(D_{\omega})}$ be a unit norm sequence that converges to $f$ in $\mathcal{H}^{1}(\varphi)$. From the continuity of $C_{\omega}$ (Lemma 18) we deduce that $C_{\omega}f_{n}\to 0$ as $n\to\infty$. Let $(c^{(n)})_{n\in{\mathbb{N}}}\subset\ell^{1}(X)$ be a sequence which satisfies $D_{\omega}c^{(n)}=f_{n}$ for each $n\in\mathbb{N}$. Then
$\|c^{(n)}\|_{\ell^{1}(X)}\lesssim\|G_{\omega}c^{(n)}\|_{\ell^{1}(X)}=\|C_{\omega}f_{n}\|_{\ell^{1}(X)}\to 0\qquad(\text{as }n\to\infty).$
This implies that
$\|f_{n}\|_{{\mathcal{H}}^{1}(\varphi)}=\|D_{\omega}c^{(n)}\|_{{\mathcal{H}}^{1}(\varphi)}\lesssim\|c^{(n)}\|_{\ell^{1}(X)}\to 0\qquad(\text{as }n\to\infty),$
which contradicts $\|f_{n}\|_{{\mathcal{H}}^{1}(\varphi)}=1$ for all $n\in\mathbb{N}$.
Finally, the last claim of Theorem 1 follows from Remark 9. ∎
5. A Comparison to ”Gabor Frames Without Inequalities”
As already mentioned in the introduction, the main insight of [21] (explicitly stated as a corollary therein) was that, for a window function $g$ in the modulation space $M^{1}({\mathbb{R}}^{d})$, the Gabor system $\mathcal{G}={\mathcal{G}}(g,\Lambda)$ over a full-rank lattice $\Lambda\subset\mathbb{R}^{2d}$ being a frame for $L^{2}(\mathbb{R}^{d})$ is equivalent to the injectivity of the analysis operator $C_{{\mathcal{G}}}:M^{\infty}({\mathbb{R}}^{d})\to\ell^{\infty}(\Lambda)$. In contrast to Theorem 1, no closed range condition is listed in this or any other of the conditions from [21, Theorem 3.1]. This is due to the fact that they are automatically fulfilled in this setting, as the next example illustrates.
Example 27.
For $g\in M^{1}(\mathbb{R}^{d})$ and $\Lambda\subset\mathbb{R}^{2d}$ a full-rank lattice, the following are equivalent:
  • (1)
    ${\mathcal{G}}={\mathcal{G}}(g,\Lambda)$ is a Gabor frame for $L^{2}(\mathbb{R}^{d})$.
  • (2)
    $C_{{\mathcal{G}}}:M^{\infty}({\mathbb{R}}^{d})\to\ell^{\infty}(\Lambda)$ is injective and has closed range.
  • (3)
    $C_{{\mathcal{G}}}:M^{\infty}({\mathbb{R}}^{d})\to\ell^{\infty}(\Lambda)$ is injective.
Indeed, (2$\Rightarrow$(3) is trivial, and (3$\Rightarrow$(1) was proven in [21]. To show (1$\Rightarrow$(2), we note that due to $g\in M^{1}(\mathbb{R}^{d})$, $S_{{\mathcal{G}}}=D_{{\mathcal{G}}}C_{{\mathcal{G}}}$ is bounded and invertible on $M^{p}(\mathbb{R}^{d})$ for all $1\leq p\leq\infty$ [21, Theorem 3.1]. From the latter, we may deduce the desired closed range condition analogously as in the proof of Theorem 1.
In a similar manner, other closed range conditions from [21, Theorem 3.1] can be derived.
At the same time, however, all closed range conditions appearing in Theorem 1 are in fact necessary. To demonstrate this, we ”naively” transfer the items (iii)-(vi), (ix), (x) and (xiii) from [21, Theorem 3.1] to our setting, obtaining the following list of conditions:
  1. $(a)$
    $S_{\psi}:{\mathcal{H}}^{\infty}(\varphi)\to{\mathcal{H}}^{\infty}(\varphi)$ is injective,
  2. $(b)$
    $C_{\psi}:{\mathcal{H}}^{\infty}(\varphi)\to\ell^{\infty}(X)$ is injective,
  3. $(c)$
    $D_{\psi}:\ell^{1}(X)\to{\mathcal{H}}^{1}(\varphi)$ has dense range,
  4. $(d)$
    $D_{\omega}:\ell^{\infty}(X)\to{\mathcal{H}}^{\infty}(\varphi)$ is injective,
  5. $(e)$
    $C_{\omega}:{\mathcal{H}}^{1}(\varphi)\to\ell^{1}(X)$ has dense range,
  6. $(f)$
    $G_{\omega}:\ell^{\infty}(X)\to\ell^{\infty}(X)$ is injective.
Below, we provide an elementary example of families $\psi$ and $\omega$ in $\mathcal{H}$, that are localized with respect to a self-localized Riesz basis $\varphi$, satisfy all the conditions (a)--(f), but are not a frame, respectively not a Riesz sequence.
Example 28.
Let $X=\mathbb{N}$, ${\mathcal{A}}$ be any solid spectral algebra, $\varphi$ be an $\mathcal{A}$-localized Riesz basis (for instance, an orthonormal basis) and define $\psi=(\psi_{k})_{k\in\mathbb{N}}:=\big{(}\frac{1}{k}\widetilde{\varphi}_{k}\big{)}_{k\in\mathbb{N}}$. Then $\omega_{k}=\frac{1}{k}S_{\varphi}^{-1/2}\varphi_{k}$ for each $k\in\mathbb{N}$, and the following hold:
  • (i)
    One has $\psi\sim_{{\mathcal{A}}}\varphi$ and, consequently, $\omega\sim_{{\mathcal{A}}}\varphi$;
  • (ii)
    $\psi$ is not a frame and, consequently, $\omega$ not a Riesz sequence in $\mathcal{H}$;
  • (iii)
    Items (a)-(f) from above are satisfied;
  • (iv)
    The range of $C_{\omega}:\mathcal{H}^{1}(\varphi)\to\ell^{1}(\mathbb{N})$ is not closed in $\ell^{1}(\mathbb{N})$.
Proof.
$(i)$ It holds $\psi\sim_{{\mathcal{A}}}\varphi$, because $G_{\psi,\varphi}=\big{[}\frac{1}{k}\delta_{k,l}\big{]}_{k,l\in\mathbb{N}}$ is contained in ${\mathcal{A}}$ due to solidity of ${\mathcal{A}}$, since $\mathcal{I}_{{\mathcal{A}}}=[\delta_{k,l}]_{k,l\in\mathbb{N}}\in{\mathcal{A}}$. The fact that $\omega\sim_{{\mathcal{A}}}\varphi$ is shown analogously or via Lemma 25.
$(ii)$ It is easy to see, that $\psi$ is not a frame, hence $\omega$ cannot be a Riesz sequence according to Theorem 1.
$(iii)$ The Gram matrix $G_{\omega}=\big{[}\frac{1}{k^{2}}\delta_{k,l}\big{]}_{k,l\in\mathbb{N}}:\ell^{\infty}(\mathbb{N})\to\ell^{\infty}(\mathbb{N})$ is obviously injective, that is, (f) is satisfied. Consequently, also (d) has to be satisfied, and (e) follows from (d) via duality upon an application of Lemma 21.
Next, we show that (a) is satisfied. Let $f=[(f_{n})_{n\in\mathbb{N}}]_{\sim_{\varphi}},g=[(g_{n})_{n\in\mathbb{N}}]_{\sim_{\varphi}}\in\mathcal{H}^{\infty}(\varphi)$ and assume that $S_{\psi}(f-g)=\sum_{l\in\mathbb{N}}\frac{1}{l^{2}}\langle f-g,\widetilde{\varphi}_{l}\rangle\widetilde{\varphi}_{l}=0$. Then $0=C_{\varphi}S_{\psi}(f-g)=\big{(}\frac{1}{k^{2}}\langle f-g,\widetilde{\varphi}_{k}\rangle\big{)}_{k\in\mathbb{N}}$. Thus $0=\langle f-g,\widetilde{\varphi}_{k}\rangle=\lim_{n\to\infty}\langle f_{n}-g_{n},\widetilde{\varphi}_{k}\rangle$ for each $k\in\mathbb{N}$ and in particular $(f_{n})_{n\in\mathbb{N}}\sim_{\varphi}(g_{n})_{n\in\mathbb{N}}$. This means that $f-g=0\in\mathcal{H}^{\infty}(\varphi)$ and (a) is shown. Statement (b) follows immediately from (a), and (c) follows from (b) via duality and Lemma 21.
$(iv)$ Since we have already shown that $C_{\omega}:\mathcal{H}^{1}(\varphi)\to\ell^{1}(\mathbb{N})$ has dense range, it suffices to show that $C_{\omega}:\mathcal{H}^{1}(\varphi)\to\ell^{1}(\mathbb{N})$ is not surjective. But the latter is clear, since $C_{\omega}=\left(\bigoplus_{k\in\mathbb{N}}\frac{1}{k}\right)C_{\varphi}S_{\varphi}^{-1/2}$, where $C_{\varphi}S_{\varphi}^{-1/2}:\mathcal{H}^{1}(\varphi)\to\ell^{1}(\mathbb{N})$ is bijective (this follows from Corollary 17), hence $\text{Ran}_{1}(C_{\omega})=\{(c_{k})_{k\in\mathbb{N}}:c_{k}=a_{k}/k,(a_{k})_{k\in\mathbb{N}}\in\ell^{1}(X)\}$ which is not equal to $\ell^{1}(X)$. ∎
Remark 29.
Via duality (Lemma 21) and the closed range theorem, we may conclude from Example 28 (iv), that $D_{\omega}:\ell^{\infty}(\mathbb{N})\to\mathcal{H}^{\infty}(\varphi)$ does not have closed range as well. Furthermore, arguing as in the proof of (5$\Leftrightarrow$(6), one may show that $D_{\omega}:\ell^{\infty}(\mathbb{N})\to\mathcal{H}^{\infty}(\varphi)$ having closed range is actually equivalent to $C_{\psi}:\mathcal{H}^{\infty}(\varphi)\to\ell^{\infty}(\mathbb{N})$ having closed range. Thus $C_{\psi}:\mathcal{H}^{\infty}(\varphi)\to\ell^{\infty}(\mathbb{N})$ does not have closed range in the above example.
Similarly, by switching the roles of $\psi$ and $\omega$ in Example 28, one readily sees that it is possible to find $\psi$ and $\omega$ which are not a frame (respectively a Riesz sequence), violate the remaining closed range conditions appearing in Theorem 1, while meeting all other conditions appearing in the same result.
Thus, each of the closed range conditions appearing in Theorem 1 is necessary for the statement to remain true.
6. Shift-Invariant Spaces
In this section, we apply Theorem 1 to the setting of shift-invariant systems. For a comprehensive overview to shift-invariant spaces, we refer the reader to [2].
Let $T_{k}$ denote the translation operator on $L^{p}(\mathbb{R}^{d})$ ($k\in\mathbb{R}^{d}$), given by $T_{k}f(t)=f(t-k)$. For a generator $\varphi\in L^{2}({\mathbb{R}}^{d})$, let $V^{p}(\varphi)$ be the shift-invariant space (formally) defined by
$V^{p}(\varphi)=\left\{\sum_{k\in{\mathbb{Z}}^{d}}c_{k}T_{k}\varphi:\ c\in\ell^{p}({\mathbb{Z}}^{d})\right\}\qquad(1\leq p\leq\infty).$
We assume that
  1. (i)
    $\varphi$ is continuous;
  2. (ii)
    $|\varphi(x)|\lesssim(1+|x|)^{-s}$ for some $s>d$; and
  3. (iii)
    $(T_{k}\varphi)_{k\in{\mathbb{Z}}^{d}}$ forms a Riesz basis for $V^{2}(\varphi)$.
Then the assumptions (i)--(iii) ensure that $V^{p}(\varphi)$ is a closed subspace of $L^{p}({\mathbb{R}}^{d})$ ($1\leq p\leq\infty$), and that $V^{2}(\varphi)$ is a reproducing kernel Hilbert space with kernel $K_{x}^{\varphi}$, i.e.
(15) $f(x)=\langle f,K_{x}^{\varphi}\rangle,\qquad f\in V^{2}(\varphi),x\in{\mathbb{R}}^{d}.$
It was shown in [19] that
$|\langle T_{k}\varphi,T_{l}\varphi\rangle|\lesssim(1+|k-l|)^{-s},\qquad k,l\in\mathbb{Z}^{d},$
i.e., the Riesz basis $(T_{k}\varphi)_{k\in{\mathbb{Z}}^{d}}$ is intrinsically ${\mathcal{J}}_{s}(\mathbb{Z}^{d})$-localized, where ${\mathcal{J}}_{s}({\mathbb{Z}}^{d})$ denotes the Jaffard algebra from Example 4 (1). Furthermore, from [19] we also know that the associated co-orbit spaces $\mathcal{H}^{p}\big{(}(T_{k}\varphi)_{k\in\mathbb{Z}^{d}}\big{)}$ coincide with the shift-invariant spaces $V^{p}(\varphi)$ ($1\leq p\leq\infty$) and
(16) $|\langle K_{x}^{\varphi},T_{l}\varphi\rangle|=|\varphi(x-l)|\lesssim(1+|x-l|)^{-s},\qquad x\in\mathbb{R}^{d},l\in\mathbb{Z}^{d}.$
Let $X=(x_{k})_{k\in{\mathbb{Z}}^{d}}\subset{\mathbb{R}}^{d}$ be discrete. If $\psi=(\psi_{k})_{k\in{\mathbb{Z}}^{d}}:=(K_{x_{k}}^{\varphi})_{k\in{\mathbb{Z}}^{d}}$ forms a frame for $V^{2}(\varphi)$ with frame bounds $0<A\leq B$, then $X$ is called a stable set of sampling for $V^{2}(\varphi)$, since, due to (15), the frame inequality (1) takes the form
$A\|f\|_{L^{2}(\mathbb{R}^{d})}^{2}\leq\sum_{k\in{\mathbb{Z}}^{d}}|f(x_{k})|^{2}\leq B\|f\|_{L^{2}(\mathbb{R}^{d})}^{2},\qquad f\in V^{2}(\varphi),$
in this case. Similarly, $X$ is called a stable set of sampling for $V^{p}(\varphi)$ ($1\leq p\leq\infty$), if there exist constants $A_{p},B_{p}>0$ such that
$A_{p}\|f\|_{L^{p}(\mathbb{R}^{d})}\leq\|(f(x_{k}))_{k\in{\mathbb{Z}}^{d}}\|_{\ell^{p}({\mathbb{Z}}^{d})}\leq B_{p}\|f\|_{L^{p}(\mathbb{R}^{d})},\qquad f\in V^{p}(\varphi).$
Now, assume that $X=(x_{k})_{k\in\mathbb{Z}^{d}}\subset{\mathbb{R}}^{d}$ is relatively separated and satisfies the mild perturbation condition
(17) $|x_{k}-k|\leq C,\qquad\text{for some }C>0\text{ and every }k\in{\mathbb{Z}}^{d}.$
Then, since
$(1+|k-l|)^{s}\leq(1+|k-x_{k}|)^{s}(1+|x_{k}-l|)^{s},\qquad k,l\in\mathbb{Z}^{d},$
the estimate (16) implies
$|\langle\psi_{k},T_{l}\varphi\rangle|\lesssim(1+|k-x_{k}|)^{s}(1+|k-l|)^{-s}\lesssim(1+|k-l|)^{-s},\qquad k,l\in\mathbb{Z}^{d},$
which means that $\psi$ and $(T_{k}\varphi)_{k\in\mathbb{Z}^{d}}$ are mutually ${\mathcal{J}}_{s}(\mathbb{Z}^{d})$-localized. Thus all assumptions of Theorem 1 are met in this setting and all of the conditions (1)--(10) provide necessary and sufficient conditions for such a set $X$ being a stable set of sampling. Instead of repeating all these conditions, we will only reformulate conditions (1), (4) and (8)--(10), since they seem more insightful to us in this case than the remaining ones.
Before doing so, we compute the R-dual $\omega$ of $\psi$. Since the frame operator $S_{\varphi}$ commutes with translations by $k\in{\mathbb{Z}}^{d}$, so does $S_{\varphi}^{-1/2}$, and we see that $S_{\varphi}^{-1/2}T_{k}\varphi=T_{k}S_{\varphi}^{-1/2}\varphi=T_{k}\gamma$ for $\gamma:=S_{\varphi}^{-1/2}\varphi$ and $k\in{\mathbb{Z}}^{d}$. Consequently,
(18) $\omega_{k}=\sum_{l\in{\mathbb{Z}}^{d}}\langle\psi_{l},T_{k}\varphi\rangle S_{\varphi}^{-{1}/{2}}T_{l}\varphi=\sum_{l\in{\mathbb{Z}}^{d}}\overline{\varphi(x_{l}-k)}T_{l}\gamma,\qquad k\in{\mathbb{Z}}^{d}.$
Note, that in case $X=(x_{k})_{k\in{\mathbb{Z}}^{d}}={\mathbb{Z}}^{d}$, the R-dual $\omega$ of $\psi$ is again a family of shifts of a single generator $\eta$ as
$\omega_{k}=T_{k}\sum_{l\in{\mathbb{Z}}^{d}}\overline{\varphi(l-k)}T_{l-k}\gamma=T_{k}\sum_{l\in{\mathbb{Z}}^{d}}\overline{\varphi(l)}T_{l}\gamma=T_{k}\eta,\qquad k\in{\mathbb{Z}}^{d}.$
Computing the entries of the associated Gram matrix $G_{\omega}$, we see that
$[G_{\omega}]_{k,n}$ $=\langle\omega_{n},\omega_{k}\rangle=\left\langle\sum_{l\in{\mathbb{Z}}^{d}}\overline{\varphi(x_{l}-n)}S_{\varphi}^{-1/2}T_{l}\varphi,\sum_{m\in{\mathbb{Z}}^{d}}\overline{\varphi(x_{m}-k)}S_{\varphi}^{-1/2}T_{m}\varphi\right\rangle$
(19) $=\sum_{l\in{\mathbb{Z}}^{d}}\varphi(x_{l}-k)\overline{\varphi(x_{l}-n)},$
since $(S_{\varphi}^{-1/2}T_{l}\varphi)_{l\in{\mathbb{Z}}^{d}}$ is an orthonormal basis.
Theorem 30.
Assume that $\varphi\in L^{2}(\mathbb{R}^{d})$ satisfies the conditions (i)--(iii), and let $(x_{k})_{k\in\mathbb{Z}^{d}}\subset\mathbb{R}^{d}$ be relatively separated. Let $G_{\omega}$ be the autocorrelation matrix with entries given by (6) and assume that condition (17) holds. Then the following are equivalent.
  • (a)
    $(x_{k})_{k\in\mathbb{Z}^{d}}$ is a stable set of sampling for $V^{2}(\varphi)$.
  • (b)
    $(x_{k})_{k\in\mathbb{Z}^{d}}$ is a stable set of sampling for $V^{\infty}(\varphi)$ and $D_{\psi}:\ell^{\infty}\to V^{\infty}(\varphi)$ has closed range.
  • (c)
    $G_{\omega}$ is invertible on $\ell^{1}({\mathbb{Z}}^{d})$.
  • (d)
    $G_{\omega}$ is invertible on $\ell^{\infty}({\mathbb{Z}}^{d})$.
  • (e)
    $G_{\omega}$ is invertible on $\ell^{2}({\mathbb{Z}}^{d})$.
Remark 31.
In the language of [26], the assumptions (i)--(iii) guarantee that $\varphi$ is contained in the Wiener amalgam space $W_{0}=W(C,\ell^{1})$ and has stable integer shifts, which, by [26, Theorem 3.1], yields several equivalent conditions for a relatively separated set $(x_{k})_{k\in{\mathbb{Z}}}\subset\mathbb{R}$ being a stable set of sampling for $V^{2}(\varphi)$. In particular, using a non-commutative version of Wiener’s lemma ([25, Proposition 8.1]), the authors of [26] showed that $(x_{k})_{k\in{\mathbb{Z}}^{d}}$ being a stable set of sampling for $V^{p}(\varphi)$ for some $1\leq p\leq\infty$ is equivalent to $(x_{k})_{k\in{\mathbb{Z}}^{d}}$ being a stable set of sampling for $V^{p}(\varphi)$ for all $1\leq p\leq\infty$ (hence the closed range condition of $D_{\psi}:\ell^{\infty}\to V^{\infty}(\varphi)$ from (b) can actually be omitted). Connected to that, we assume condition (17) in order to guarantee localization with respect to the Jaffard algebra ${\mathcal{J}}_{s}(\mathbb{Z}^{d})$ and utilize inverse-closedness in ${\mathcal{B}}(\ell^{2}(\mathbb{Z}^{d}))$, which can be viewed as another non-commutative version of Wiener’s lemma.
We also note that the autocorrelation matrix $G_{\omega}$ is given by the product $G_{\omega}=P_{X}(\varphi)^{*}P_{X}(\varphi)$, where $P_{X}(\varphi)=(\varphi(x_{l}-k))_{l,k\in{\mathbb{Z}}}$ is the matrix appearing in [26, Theorem 3.1]. According to [26, Theorem 3.1] (in the case $d=1$), stable sampling is equivalent to $P_{\Gamma}(\varphi):\ell^{\infty}(\mathbb{Z}^{d})\to\ell^{\infty}(\Gamma)$ being bounded and injective for all weak limits $\Gamma$ of integer translates of $(x_{k})_{k\in{\mathbb{Z}}}$. In contrast, conditions (c)--(e) from above only concern one specific operator $G_{\omega}$. To the best of our knowledge, these invertibility conditions on the autocorrelation matrix $G_{\omega}$ are new.
Acknowledgements
The authors would like to thank K. Gröchenig for pointing out the connection between Theorem 30 and [26].
This research was funded by the Austrian Science Fund (FWF) 10.55776/P34624 (P.B. and L.K.) and 10.55776/PAT1384824 (M.S.). For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.